### v1

# N-GLARE: An Non-Generative Latent Representation-Efficient LLM Safety Evaluator

Abstract
Evaluating the safety robustness of LLMs is critical for their deployment. However, mainstream Red Teaming methods rely on online generation and black-box output analysis. These approaches are not only costly but also suffer from feedback latency, making them unsuitable for agile diagnostics after training a new model. To address this, we propose N-GLARE (A Non-Generative, Latent Representation-Efficient LLM Safety Evaluator). N-GLARE operates entirely on the model’s latent representations, bypassing the need for full text generation. It characterizes hidden layer dynamics by analyzing the APT (Angular-Probabilistic Trajectory) of latent representations and introducing the JSS (Jensen-Shannon Separability) metric. Experiments on over 40 models and 20 red teaming strategies demonstrate that the JSS metric exhibits high consistency with the safety rankings derived from Red Teaming. N-GLARE reproduces the discriminative trends of large-scale red-teaming tests at less than 1% of the token cost and the runtime cost, providing an efficient output-free evaluation proxy for real-time diagnostics.
1 Introduction
In recent years, Large Language Models (LLMs) have developed rapidly, with their capability boundaries continuously expanding. They have been deployed in a wide range of application scenarios, exerting substantial impact on social production and daily life [grattafiori2024llama, zhou2024large]. However, alongside this growth in capability, security risks and alignment issues have become increasingly prominent, emerging as a critical bottleneck for large-scale, safety-critical deployment [das2025security, dong2024attacks]. Trained on vast and often uncurated Internet corpora, LLMs can inadvertently internalize social biases and potentially harmful knowledge. Without robust safety alignment, they may produce unsafe outputs when exposed to malicious prompts, complex contexts, or specific triggers—for example, generating cyberattack code [wei2023jailbroken], spreading misleading information [russinovich2025great], or guiding embodied agents toward physically harmful behaviors [yang2025concept]. Ensuring that LLMs remain aligned with human values and safety norms is therefore a central challenge.
Evaluating model safety is a foundational step toward achieving such alignment. The dominant paradigm in both industry and academia is currently Red Teaming [lin2025against], which simulates adversarial user behavior by designing diverse, subtle attack prompts to probe the model and elicit unsafe outputs. While this online, behavior-based testing has proved effective for discovering vulnerabilities, it suffers from two fundamental limitations. First, it is highly resource-intensive: test coverage depends on the scale and diversity of attack prompts, which require continual human and computational investment as new attack vectors emerge. Second, Red Teaming is essentially a black-box methodology. It only observes input–output behavior and provides little insight into the model’s internal safety state. A refusal that looks safe at the surface level may simply be triggered by brittle pattern matching, rather than reflecting a robust internal representation of harmfulness.
To move beyond purely behavioral stress testing, recent work has begun to analyze LLMs in their latent representation space [zou2023representation, li2025revisiting]. Representation engineering shows that when a model processes different concepts (including abstract notions such as “truthfulness,” “refusal,” or “harmfulness”), its hidden activations form structured, often linearly separable patterns. Let $h_{l,p}\in\mathbb{R}^{d}$ denote the activation at layer $l$ and position $p$. For two contrasting semantic sets $\mathcal{S}^{+}$ and $\mathcal{S}^{-}$, their mean activations differ along concept directions of the form
$o\approx\mathbb{E}[H(\mathcal{S}^{+})]-\mathbb{E}[H(\mathcal{S}^{-})],$ (1)
suggesting that high-level semantics are explicitly encoded in latent geometry. Existing work primarily uses such directions at the prompt level, for example to steer individual generations or to classify the risk of a specific input. However, these techniques are typically model-specific and do not directly provide a model-level safety indicator that is comparable across architectures and checkpoints.
In this paper, we take a step in that direction and ask: can we evaluate model safety without generating full outputs, solely by examining how its internal representations evolve under different safety conditions? To this end, we propose N-GLARE (A Non-Generative, Latent Representation-Efficient LLM Safety Evaluator), a framework that operates entirely in latent space and bypasses token-level generation. For each model, we construct Angular-Probabilistic Trajectories (APT) of hidden representations under different probing conditions: benign ($B$), jailbreak ($J$), plainquery ($P$) and an idealized refusal counterfactual ($R$). After normalizing these trajectories along a standardized progress axis, we model benign activations as a low-dimensional manifold and define a geometric turning angle that measures how strongly a trajectory deviates away from this benign manifold over time. We then introduce Jensen–Shannon Separability (JSS), which compares the slice-wise distributions of these turning angles across conditions, and aggregates them into model-level safety indicators, introducing other JSS-based safety scores.
Figure 1: JSS Shows Stronger Safety Alignment in Qwen3-4B-SafeRL Compared to Base and Alignment-Removed Variants. Statistics on the JSS distribution for 100 jailbreak trajectories and the ideal refusal pattern.
Figure 1 illustrates the core intuition: when comparing multiple variants of the same base model (e.g., RL-aligned, base, and alignment-removed versions of Qwen3-4B), we observe that better-aligned variants exhibit more pronounced geometric separation between benign and jailbreak trajectories in latent space, and their jailbreak trajectories are also further away from the ideal refusal pattern. This suggests that internal trajectory separability can serve as a proxy for the underlying safety state of the model, potentially reproducing the discriminative power of large-scale Red Teaming at a fraction of the cost.
We formalize this idea in Section 3 by casting safety evaluation as a problem of comparing representation trajectories under $(B,J,R,P)$ conditions. We then instantiate our method in Section 4: benign activations define a whitened manifold, the trajectory turning angle encodes instantaneous deviation from this manifold, and Jensen–Shannon divergence over angle distributions along the progress axis yields JSS metrics that are both geometrically interpretable and amenable to cross-model comparison. Experiments on over 40 models and 20 red-teaming strategies demonstrate that N-GLARE faithfully reproduces the safety rankings induced by traditional Red Teaming while using less than 1% of the token and runtime cost, making it suitable for agile diagnostics after training new models or updating safety mechanisms.
Our main contributions are summarized as follows:
  • A non-generative, latent-space safety evaluator. We propose N-GLARE, a framework that entirely bypasses online text generation and external output evaluators. To our knowledge, this is the first work to leverage latent representation trajectories for model-level LLM safety assessment.
  • A geometric formalization of safety alignment. We formulate safety evaluation as a trajectory separability problem in representation space, construct a benign manifold via whitened PCA, and define a introduce Angular-Probabilistic Trajectories (APT) deviation measure. Building on this, we propose Jensen–Shannon Separability (JSS) and JSS-related normalized safety score as statistically robust, cross-model comparable indicators.
  • Extensive empirical validation and efficiency gains. Across more than 40 models and 20 red-teaming strategies, we show that JSS-based rankings are highly consistent with those derived from traditional Red Teaming, while reducing token and runtime cost to below 1%. We further demonstrate that N-GLARE supports fine-grained analyses such as early-stage safety dynamics and comparison of alignment interventions across checkpoints, providing a practical tool for real-time safety diagnostics.
2 Technical Background
2.1 LLM Security and Red Teaming
Despite the widespread implementation of safety alignment in contemporary mainstream LLMs, they continue to face extensive security challenges in practical application [Cui2024Risk]. Attackers frequently employ meticulously designed adversarial inputs to bypass the model’s safety guardrails [zhu2023promptbench, paulus2024advprompter]. Through so-called “jailbreak attacks,” models are induced to circumvent their established policies, consequently generating harmful content. Such content may be exploited for academic misconduct [oliveira2025human], facilitating cyberattacks [xu2024autoattacker], or infringing upon copyright [xu2024llms].
Consequently, ensuring that model outputs adhere to ethical norms and effectively mitigating potential risks have become indispensable research directions in the field of artificial intelligence. In this context, the security evaluation of large models is paramount, wherein Red Teaming has emerged as a core methodology for assessing model safety.
Red Teaming is an adversarial evaluation method centered on proactively discovering and identifying potential model vulnerabilities by simulating attacker strategies and behaviors. Existing Red Teaming approaches can be broadly categorized into two types. The first focuses on content safety in single-turn interactions, such as leveraging algorithms like Monte Carlo Tree Search (MCTS) and evolutionary algorithms to optimize attack prompts [sadasivan2024fast], or generating jailbreak templates through the mutation of seed prompts [yu2023gptfuzzer]. The second category addresses security risks in multi-turn dialogue scenarios, typically employing specialized agents for attack planning, executing dynamic multi-turn jailbreaks, and evaluating the attack’s efficacy [rahman2025x, weng2025foot, ren2025llms].
Although existing methods show a trend towards diversification, their mainstream paradigm still generally follows a three-stage process: first, constructing the attack prompt; second, obtaining the target model’s output for the prompt; and finally, judging the success of the attack based on the evaluation of the model’s output results. This model leads to particularly high evaluation overhead for Red Teaming.
2.2 LLM Latent Space Analysis Techniques
The internal mechanisms of LLMs are often treated as a black box. To understand how they process information, researchers analyze their latent space, i.e., the internal high-dimensional hidden states. In Transformers, this corresponds to the activation vector $h_{l,p}\in\mathbb{R}^{d}$ at a specific layer $l$ and token position $p$, where $d$ is the hidden dimension.
2.2.1 Embedding Analysis
Embedding analysis reveals semantic meaning by examining the static geometry of latent space vectors. Common methods like cosine similarity ($Sim(\vec{A},\vec{B})=\frac{\vec{A}\cdot\vec{B}}{\|\vec{A}\|\|\vec{B}\|}$) and dimensionality reduction visualization like t-SNE can reveal semantic clusters (e.g., “king” with “queen”) [rogers2020primer]. Furthermore, domain-adaptation analysis shows that fine-tuning (e.g., BioBERT [lee2020biobert]) reshapes the latent space, forming more refined domain-specific subspaces to enhance downstream task performance.
2.2.2 Probing and Intervention
This technique shifts from passive analysis to active decoding and control. Early probing [tenney2019bert] used simple linear classifiers ($y=f(H)$) to test if latent activations $H$ encoded specific linguistic properties, discovering that information is stored hierarchically in LLMs.
A more advanced paradigm, Representation Engineering (RepE) [zou2023representation], controls model behavior by directly manipulating latent representations. Its core is identifying vector directions, or concept operators $o^{c}$, that represent high-level concepts (e.g., truthfulness). This is often achieved by contrasting the activation differences from different input sets ($\mathcal{S}^{c+}$ vs $\mathcal{S}^{c-}$):
$o^{c}\approx\mathbb{E}[H(\mathcal{S}^{c+})]-\mathbb{E}[H(\mathcal{S}^{c-})]$ (2)
Here, $H(\mathcal{S})$ represents the mean activation of model $\mathcal{M}$ when processing a specific input set $\mathcal{S}$. RepE is widely applied in LLM safety [cai2024self, wang2025model, wang2024inferaligner, cao2024nothing], for instance, by using positive or negative steering vectors to enhance truthfulness [marks2023geometry, li2023inference, wang2025adaptive, rimsky2024steering] or suppress meta-behaviors like “refusal” [arditi2024refusal].
2.2.3 Attention Mechanism Analysis
Attention mechanism analysis interprets the model’s information flow via its weight distribution $A=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)$. Here, the $Q,K,V$ matrices are linear projections of the layer input $h_{l-1}$. Heatmap visualization [vig2019multiscale] reveals that attention exhibits hierarchical properties (from syntax to semantics across layers). At the same time, research [michel2019sixteen] indicates high redundancy in attention mechanisms. To address its computational complexity, models like Perceiver [jaegle2021perceiver] introduce low-dimensional latent variables to compress the computation.
2.2.4 Latent Space Dynamic Analysis
Dynamic analysis focuses on the activation trajectories during the reasoning process. Researchers not only track attention changes during Chain-of-Thought (CoT) [wei2022chain] but also begin to analyze the latent space trajectories themselves. For example, Li et al. [li2025core] proposed Chain-of-Reasoning Embedding (CoRE), which treats the reasoning process as a series of hidden states $h_{l,p_{T}}$. They found that redundant reasoning (”overthinking”) manifests as identifiable cyclical fluctuations in the trajectory. Based on this, the CoRE-Eval framework was developed, enabling label-free self-evaluation to dynamically terminate inefficient reasoning, thereby improving efficiency.
Despite significant progress in analysis techniques for LLM latent spaces, the field still faces the challenge of model-specificity: the identified features, concept operators (like $o^{c}$), attention heads, or reasoning trajectories are often highly dependent on a single model $\mathcal{M}$ and lack transferability across models.
3 Problem Statement
Safety evaluation of large language models is currently dominated by generative red-teaming: a model is queried with adversarial prompts generated by red-teaming LLMs, asked to produce multi-token outputs, and then judged by a human or a separate LLM-based evaluator. This approach is expensive (because evaluation cost grows with generated tokens), noisy (because harmfulness must be inferred from surface text), and incomplete (because it reveals only the final behavior, not the underlying safety state).
Formally, let $M_{\text{subject}},M_{s}$ be a model and let $x$ be a potentially harmful input drawn from an adversarial distribution $\mathcal{D}_{\mathrm{adv}}$ or online $M_{\text{attacker}},M_{r}$. A generative evaluator, another LLM $M_{\text{evaluator}},M_{e}$ observes the decoded output $y=M_{s}(x)$ and produces an assessment $\mathrm{Eval_{M_{e}}}(y)$ that approximates the latent safety function $f_{\mathrm{safety}}(M_{s},x)$. That is,
$\mathrm{Eval_{M_{e}}}(y)\approx f_{\mathrm{safety}}(M_{s},x),$
where this 0/1 function, $f_{\mathrm{safety}}$ represents whether the model processes $x$ in a safe or unsafe manner. This evaluation pipeline is entirely output-dependent (by analyzing the natural language text) and therefore cannot access $f_{\mathrm{safety}}(M_{s},x)$ directly.
In contrast, a non-generative evaluator also seeks to estimate $f_{\mathrm{safety}}(M_{s},x)$ without decoding any output tokens. Instead, it relies only on information available inside the $M_{s}$. The goal is to design a mapping
$g:(M_{s},x)\;\mapsto\;f_{\text{safety}}(M_{s},x),$
where $g(M,x)$ is a model-level safety signal that: (i) approximates the same generative-based quantity $\mathrm{Eval_{M_{e}}}(y)$ evaluated by LLM-based evaluator, (ii) is comparable across models and prompts, and (iii) does not rely on online red-teaming text generation or attackers $M_{r}$.
The core problem addressed in this work is therefore:
How can we construct a non-generative safety metric $g(\cdot,\cdot)$ that reliably reflects a model’s internal safety state and correlates with its behavioral robustness observed under traditional red-teaming?
This paper develops such a metric by operating entirely in the representation space, enabling low-cost, model-level safety evaluation without generation.
4 Methodology
We now instantiate the representation-space safety evaluation framework outlined above. Given a set of models $\mathcal{M}=\{M_{1},\dots,M_{K}\}$ and probing conditions $A\in\{B,J,R,P\}$ (benign, jailbreak, ideal refusal), our goal is to map internal activation trajectories to scalar safety indicators. The pipeline proceeds in three stages: (i) extracting layer-wise hidden states and normalizing them along a standardized progress axis; (ii) fitting a benign manifold and defining a geometric turning angle that measures instantaneous deviation from this manifold; and (iii) aggregating slice-wise Jensen–Shannon divergences of turning-angle distributions into Jensen–Shannon Separability (JSS) metrics and a normalized safety score.
4.1 Trajectory Extraction and Standardized Progress
For each model $M$ and probing condition $A\in\{B,J,R,P\}$ (defined in Section 4.3), we construct probing sequences that form either genuine interaction histories (e.g., multi-turn dialogues) or canonical prefix expansions for single-turn prompts. At each node $t$ in such a sequence, we identify a fixed decision-making cross-section, chosen in our experiments as the point immediately before the model generates the next assistant response. A single forward pass at node $t$ yields the last-token hidden activations across all transformer layers,
$\Phi_{M}^{(A)}(t)=\bigl(h_{L,t}^{(A)},\dots,h_{1,t}^{(A)}\bigr),\quad h_{\ell,t}^{(A)}\in\mathbb{R}^{d_{\ell}},$ (3)
where $L$ is the number of layers and $d_{\ell}$ is the hidden dimension at layer $\ell$.
To improve cross-model comparability and reduce dimensionality, we partition layers into a small number of layer groups $G\in\{\text{lower},\text{middle},\text{upper}\}$ and average within each group, obtaining group-level representations
$h_{t}^{(A,G)}\in\mathbb{R}^{d},$ (4)
with a common dimension $d$. For each condition $A$ and group $G$, the sequence $\{h_{t}^{(A,G)}\}_{t=1}^{T_{A}}$ defines a trajectory in latent space.
The length $T_{A}$ of a probing sequence may vary across prompts and models. To enable comparison on a common timeline, we normalize each trajectory by its arc length and define a standardized progress variable $s_{t}\in[0,1]$:
$s_{1}=0,\qquad s_{t}=\frac{\sum_{k=1}^{t-1}\|h_{k+1}^{(A,G)}-h_{k}^{(A,G)}\|_{2}}{\sum_{k=1}^{T_{A}-1}\|h_{k+1}^{(A,G)}-h_{k}^{(A,G)}\|_{2}},\quad t>1.$ (5)
This parametrization treats trajectories from different prompts and models as curves evolving along the shared progress axis $s\in[0,1]$. In single-turn settings where there is no natural temporal dimension, we construct pseudo-sequences via canonical prefix augmentation (e.g., from an empty system prompt to the full user query) and apply the same normalization. We denote the resulting standardized trajectory under condition $A$ and group $G$ by
$\Gamma_{M}^{(A,G)}=\{(h_{t}^{(A,G)},s_{t})\}_{t=1}^{T_{A}}.$ (6)
4.2 Benign Manifold and Geometric Turning Angle
To quantify deviations from normal behavior, we first construct a benign manifold using representations collected under the benign condition $B$. For each group $G$, we aggregate all benign activations across prompts and progress values:
$\mathcal{H}_{B}^{(G)}=\{h^{(B,G)}\}.$ (7)
We estimate the empirical mean and covariance
$\begin{split}\mu^{(G)}&=\mathbb{E}[h^{(B,G)}],\qquad\\ \Sigma^{(G)}&=\mathbb{E}\bigl[(h^{(B,G)}-\mu^{(G)})(h^{(B,G)}-\mu^{(G)})^{\top}\bigr],\end{split}$ (8)
and perform a rank-$r$ eigendecomposition
$\Sigma^{(G)}\approx U_{r}^{(G)}\Lambda_{r}^{(G)}{U_{r}^{(G)}}^{\top},\quad r\ll d,$ (9)
where $U_{r}^{(G)}\in\mathbb{R}^{d\times r}$ contains the top-$r$ eigenvectors and $\Lambda_{r}^{(G)}\in\mathbb{R}^{r\times r}$ is the diagonal matrix of corresponding eigenvalues. This defines a benign subspace
$\mathcal{S}^{(G)}=\mathrm{span}\bigl(U_{r}^{(G)}\bigr),$ (10)
and a whitening transform
$W^{(G)}={\Lambda_{r}^{(G)}}^{-1/2}{U_{r}^{(G)}}^{\top}.$ (11)
For any representation $h$, its whitened coordinate is
$z=W^{(G)}(h-\mu^{(G)}),$ (12)
which removes scale and correlation effects in the benign distribution and provides a canonical coordinate system for measuring deviations across models.
By the Eckart–Young–Mirsky theorem [EckartYoung1936, GolubVanLoan2013] and the tangent-space linearization principle for smooth manifolds [DoCarmo1976], the top-$r$ PCA subspace offers the statistically optimal linear surrogate of the benign manifold.
Given the benign manifold, we define an instantaneous geometric measure that characterizes how a trajectory moves relative to this manifold. For any representation $h$, we consider its reconstruction from the benign subspace:
$\hat{h}^{(G)}=\mu^{(G)}+U_{r}^{(G)}{U_{r}^{(G)}}^{\top}(h-\mu^{(G)}),$ (13)
and define the residual
$r^{(G)}(h)=h-\hat{h}^{(G)}=h-\mu^{(G)}-U_{r}^{(G)}{U_{r}^{(G)}}^{\top}(h-\mu^{(G)}).$ (14)
The squared norm
$E_{b}^{(G)}(h)=\|r^{(G)}(h)\|_{2}^{2}$ (15)
acts as a deviation energy from the benign manifold. Its gradient
$n^{(G)}(h)=\nabla E_{b}^{(G)}(h)=2r^{(G)}(h)$ (16)
defines an outward normal direction pointing away from the benign subspace at $h$.
For each trajectory and its representation $h_{t}\in\Gamma_{M}^{(A,G)}$, the tangent direction at node $t$ is given by
$\tau_{t}^{(A,G)}=\frac{h_{t+1}^{(A,G)}-h_{t}^{(A,G)}}{\|h_{t+1}^{(A,G)}-h_{t}^{(A,G)}\|_{2}},$ (17)
for $1\leq t<T_{A}$. We then define the turning angle $\theta_{t}^{(A,G)}\in[0,\pi]$ between the trajectory tangent and the outward normal:
$\theta_{t}^{(A,G)}=\arccos\left(\frac{\tau_{t}^{(A,G)\top}n^{(G)}\bigl(h_{t}^{(A,G)}\bigr)}{\|\tau_{t}^{(A,G)}\|_{2}\,\|n^{(G)}\bigl(h_{t}^{(A,G)}\bigr)\|_{2}}\right).$ (18)
In practice, since $n^{(G)}(h)$ is collinear with $r^{(G)}(h)$, we compute the numerator as $\tau_{t}^{(A,G)\top}r^{(G)}\bigl(h_{t}^{(A,G)}\bigr)$ and normalize accordingly. Intuitively, $\theta_{t}^{(A,G)}$ measures how strongly the trajectory at $t$ turns “away” from the benign manifold. Smaller angles (particularly $\theta_{t}^{(A,G)}<\pi/2$) indicate movement into regions further from benign behavior.
Algorithm 1 JSS Core Computation for Model $M$
0: Model $M$; datasets $\mathcal{D}_{A}$ for $A\in\{B,J,R,P\}$; hyperparameters $(r,G,I,s_{0},\{w_{i}\},\{w_{i}^{\prime}\})$.
[…]
1:Trajectory extraction and benign manifold fitting
2:for each $A\in\{B,J,R,P\}$ do
3:  $\{\Gamma_{M}^{(A,G)}\}_{G}\leftarrow\textsc{ExtractAndNormalize}(\mathcal{D}_{A},M,G)$
4:end for
[…]
7:Turning-angle computation
8:for each $A\in\{B,J,R,P\}$ and each $G$ do
9:  $\Theta^{(A,G)}\leftarrow\textsc{CalculateAngles}(\Gamma_{M}^{(A,G)},\mu^{(G)},U_{r}^{(G)})$ {all $(\theta_{t}^{(A,G)},s_{t})$}
10:end for
11:Slice-wise JS aggregation
12:for each comparison $(A,B^{\prime})\in\{(J,B),(J,R),(B,R),(P,B)\}$ do
13:  for each $G$ do
14:   $\mathrm{JSS}^{(G)}(A,B^{\prime})\leftarrow 0$, $\mathrm{JSS}^{(G)}_{\text{early}}(A,B^{\prime})\leftarrow 0$
15:   for $i=1$ to II do
16:    $\Theta_{A}^{(G)}(s_{i})\leftarrow\{\theta:(\theta,s)\in\Theta^{(A,G)},s\in s_{i}\}$
17:    $\Theta_{B^{\prime}}^{(G)}(s_{i})\leftarrow\{\theta:(\theta,s)\in\Theta^{(B^{\prime},G)},s\in s_{i}\}$
18:    $\mathrm{JS}_{i}^{(G)}(A,B^{\prime})\leftarrow\mathrm{JS}\bigl(\mathrm{Dist}(\Theta_{A}^{(G)}(s_{i}))\,\big\|\,\mathrm{Dist}(\Theta_{B^{\prime}}^{(G)}(s_{i}))\bigr)$
19:    $\mathrm{JSS}^{(G)}(A,B^{\prime})\mathrel{+}=\mathrm{JS}_{i}^{(G)}(A,B^{\prime})$
20:    if $s_{i}\leq s_{0}$ then
21:     $\mathrm{JSS}^{(G)}_{\text{early}}(A,B^{\prime})\mathrel{+}=\mathrm{JS}_{i}^{(G)}(A,B^{\prime})$
22:    end if
23:   end for
24:  end for
25:  $\mathrm{JSS}(A,B^{\prime})\leftarrow\frac{1}{|G|}\sum_{G}\mathrm{JSS}^{(G)}(A,B^{\prime})$
26:  $\mathrm{JSS}_{\text{early}}(A,B^{\prime})\leftarrow\frac{1}{|G|}\sum_{G}\mathrm{JSS}^{(G)}_{\text{early}}(A,B^{\prime})$
27:end for
28:return $\mathrm{JSS}(J,B)$, $\mathrm{JSS}_{\text{early}}(J,B)$, $\mathrm{JSS}(J,R)$, $\mathrm{JSS}_{\text{early}}(J,R)$, $\mathrm{JSS}(P,B)$
4.3 Probing Conditions, Slicing, and JSS Metrics
We apply the above constructions to three probing conditions that are designed to be isomorphic along the progress axis $s$:
  • Benign ($B$): natural or conventional task distributions, used to fit the benign manifold and serve as a reference.
  • Jailbreak ($J$): adversarial sequences designed to elicit harmful behavior.
  • Ideal refusal ($R$): sequences identical to $J$ in user content, but with the assistant response replaced by a fixed, policy-compliant refusal template.
  • Plain Query ($P$): sequences that repeatedly present the raw jailbreak true intention to the model, capturing the model’s unconstrained internal response trajectory under direct exposure to harmful prompts.
For each group $G$ and condition $A$, we thus obtain a set of angle–progress pairs (i.e. Angular-Probabilistic Trajectories, APT)
$\Theta^{(A,G)}=\{(\theta_{t}^{(A,G)},s_{t})\}.$ (19)
To study how separability evolves along the interaction, we discretize the standardized axis $s\in[0,1]$ into $I$ slices (bins) $\{s_{i}\}_{i=1}^{I}$. In slice $s_{i}$, we collect all angles whose progress falls into that interval:
$\Theta_{A}^{(G)}(s_{i})=\{\theta_{t}^{(A,G)}:s_{t}\in s_{i}\}.$ (20)
We also define an early-stage threshold $s_{0}\in(0,1)$ (default $s_{0}=0.4$) and use it to construct “early-ss” metrics that focus on whether the model exhibits safe behavior in the initial portion of the trajectory.
Within each slice $s_{i}$ and group $G$, we estimate the empirical distributions of the angle samples for two conditions $A$ and $B$ (e.g., $J$ vs. $B$, or $J$ vs. $R$), denoted by $\mathrm{Dist}(\Theta_{A}^{(G)}(s_{i}))$ and $\mathrm{Dist}(\Theta_{B}^{(G)}(s_{i}))$, meaninig the distribution of them. We then compute the Jensen–Shannon divergence:
$\begin{split}\mathrm{JS}_{i}^{(G)}(A,B)=\mathrm{JS}\bigl(\mathrm{Dist}(\Theta_{A}^{(G)}(s_{i}))\,\big\|\,\mathrm{Dist}(\Theta_{B}^{(G)}(s_{i}))\bigr)\end{split}$ (21)
To obtain a model-level separability measure, we aggregate slice-level divergences. For each group $G$,
$\mathrm{JSS}^{(G)}(A,B)=\sum_{i=1}^{I}\,\mathrm{JS}_{i}^{(G)}(A,B),$ (22)
and average across layer groups to obtain the all-s separability:
$\mathrm{JSS}(A,B)=\frac{1}{|G|}\sum_{G}\mathrm{JSS}^{(G)}(A,B).$ (23)
Restricting to slices with $i\leq s_{0}$ yields the early-s separability:
$\mathrm{JSS}_{\text{early}}(A,B)=\frac{1}{|G|}\sum_{G}\sum_{i\leq s_{0}}\mathrm{JS}_{i}^{(G)}(A,B).$ (24)
Algorithm 1 summarizes the core computation of JSS metrics for a single model $M$. For clarity, the pseudocode illustrates a single run on the full dataset. In our experiments, we wrap this procedure in a bootstrap loop to obtain empirical confidence intervals and perform non-parametric statistical tests across models.
(a) Lower Group
(b) Middle Group
(c) Upper Group
(d) Lower Group
(e) Middle Group
(f) Upper Group
Figure 2: APT trajectories collected from different layer groups. Subfigures (a)(b)(c) correspond to safety-aligned models, while (d)(e)(f) are from jailbreak-tuned models. Red curves represent Plain Query trajectories, yellow denote Benign trajectories, blue indicate Jailbreak trajectories, and green correspond to Ideal Refusal trajectories. Safety-aligned LLMs exhibit well-separated and low-variance APT trajectories, whereas jailbreak-tuned LLMs show weakened separability and substantially higher variance across all query types, indicating degraded safety boundaries.
5 Experiments and Analysis
This section begins by presenting the intuitive safety characteristics derived from latent representation trajectories, illustrating how model representations evolve under different dialogue families. Subsequently, from a dynamical perspective, we systematically examine the correspondence between these latent features and the model’s surface-level safety behaviors, thereby validating the safety representation hypothesis proposed earlier. On this basis, we employ quantified safety feature metrics to evaluate various aligned models and compare their rankings with those obtained from systematic traditional red-teaming evaluations, demonstrating a high degree of consistency in safety scales as well as a significant advantage in evaluation cost. Finally, we investigate the statistical consistency of the ranking results across different datasets and sample sizes, analyzing the stability of this evaluation framework under changes in data distribution and sampling conditions.
5.1 Trajectory Evolution Dynamics
This subsection empirically establishes one core observation of our evaluation framework:
Observation 1.
A model’s safety state is not only a behavior-based static attribute, but rather a dynamic process that can be continuously tracked within the latent representation space.
Specifically, in Figure 2, we compare the APT evolutionary trajectories of the PlainQuery, Baseline, Jailbreak, and Ideal Refusal dialogue families. Baseline maintains a stable, flat trajectory, representing the model’s consistent generation direction during natural dialogue. PlainQuery rises sharply across all layers, rapidly diverging from the baseline, indicating that the model promptly shifts its generation direction upon detecting harmful intent, forming a distinct refusal response. Jailbreak trajectories gradually diverge upward from the baseline, exhibiting a notable downward drift in the higher layers, suggesting that adversarial semantics accumulate in upper-layer representations, progressively eroding the refusal mechanism. Although the Ideal-Refusal trajectory initially follows the Jailbreak path, a clear bifurcation emerges, indicating that the model encodes a latent distinction between accepting and refusing harmful instructions. Such separation suggests that the model possesses an intrinsic cognitive bias that differentiates unsafe compliance from defensive rejection within its internal representation space.
(a) Qwen Family
(b) Gemma Family
(c) DeepSeek Family
(d) LessEq 3B Family
(e) Generic 7B Family
(f) GreatEq 13B Family
Figure 3: Multi-perspective cross-model safety ranking consistency between N-GLARE and traditional red-teaming. For each model family ((a–f; see Appendix Table II for model abbreviations on scatters)), the left scatter plot compares the relative ranks induced by our JSS-based mixed proxy (X-axis) and by Hybrid red-teaming scores (Y-axis); points lying near the diagonal blue band indicate close agreement between the two rankings. The right panel in each family shows the corresponding raw JSS-proxy (blue) and Hybrid (red) scores over all subset models, with shaded regions denoting bootstrap variability, again yielding almost identical within-family orderings. Note that ”Hybrid” is the overall average across score all tested red-teaming tasks, with detailed task-wise scores provided in Appendix Table LABEL:tab:jb_safety_multi.
5.2 Synchronous Evolution of Jailbreak Progression, Refusal Tendency, and JSS Dynamics
After obtaining the intuitive empirical phenomena, we focus on verifying two key observations:
Observation 2.
Divergence between Jailbreak and Baseline trajectories as an indicator of the model’s discriminative capability to detect jailbreak attacks process shifting.
A separation between the jailbreak attack trajectory and the baseline behavior trajectory in the representation space signifies a clear process shift in the model’s internal dynamics from normal generation to adversarial alignment. This divergence reflects the model’s discriminative capability to detect and respond to emerging jailbreak cues, serving as a direct and quantifiable indicator of its latent safety awareness.
Observation 3.
Distance between Jailbreak and Ideal Refusal trajectories as an indicator of the model’s intrinsic refusal tendency.
The JS distance between Jailbreak and Ideal Refusal trajectories directly reflects the model’s ability to distinguish and resist jailbreak behaviors. An increase in this distance indicates that the model has a stronger latent discrimination between “accepting” and “blocking” jailbreak instructions, thus exhibiting an enhanced intrinsic refusal tendency, corresponding to a higher level of safety awareness.
5.2.1 Phase-wise Dynamics of Jailbreak Trajectories and Latent Safety Awareness
For observation 2, if a model lacked latent safety awareness, JB activations would remain near the Baseline manifold, yielding no meaningful temporal segmentation. To test whether phase structure exists and whether it is safety-dependent, we use a lightweight neural classifier as an operational probe: if representations at different jailbreak stages encode distinct information, the classifier should reliably separate them. Aligned Llama-family models indeed show strong separability: 0.86–0.97 (start vs. mid), 0.91–0.96 (start vs. end), 0.85–0.86 (mid vs. end). However, after ablating safety alignment (manually jailbroken variants), stage classification collapses to $\sim 0.33$ for all pairs. This contrast provides two conclusions: (1) the phase structure is statistically reliable, because it can be consistently decoded by a simple classifier; (2) it is safety-dependent, because the structure disappears once alignment is destroyed. Thus, aligned models internally encode a latent variable corresponding to the stage of the jailbreak process, and the resulting phase-wise divergence of JB from Baseline offers a direct, graded indicator of latent safety awareness.
Figure 4: Coupled dynamics of JSS and ANM along the jailbreak process. The color indicates the normalized refusal-intensity score (ANM); The inset line plot in the upper-right overlays slice-wise JSS and ANM curves over the standardized progress axis, revealing a shared mid-phase rise and late-phase collapse, which highlights the strong correlation between geometric separability (JSS) and probability-level refusal tendency (ANM).
5.2.2 Distance between Jailbreak and Ideal Refusal Trajectories and the Model’s Refusal Tendency
For another observation 3, empirical measurements confirm a strong synchrony between Ideal Refusal-Jailbreak separability and model’s refusal tendency. We evaluate the model’s latent refusal inclination strength by using a unified metric of how strongly the model tends to generate reject-related tokens at each stage of the jailbreak process. Specifically, we formalize ANM, which quantifies probability-level concentration on “refusal-cluster” tokens.
$\mathrm{ANM}(x)=\big\langle W_{\mathrm{ref}},\,p(\cdot\mid x)\big\rangle=\sum_{v\in\mathcal{V}}W_{\mathrm{ref}}(v)\,p(v\mid x),$ (25)
where $p(v\mid x)=\mathrm{softmax}(z_{v}(x)/\tau_{\text{pred}})$ is the model’s next-token distribution, and $W_{\mathrm{ref}}$ is a fixed refusal–prototype distribution constructed by:
$W_{\mathrm{ref}}(v)=\frac{\exp(\langle\hat{e}_{v},\hat{c}\rangle/\tau_{\text{ref}})}{\sum_{u\in\mathcal{V}}\exp(\langle\hat{e}_{u},\hat{c}\rangle/\tau_{\text{ref}})},$ (26)
with $\hat{c}$ obtained by averaging normalized embeddings of several refusal-seed phrases. (”sorry”, ”I can’t”, etc.) Thus, ANM quantifies how strongly the model’s predictive distribution aligns with refusal-semantics in latent space.
Across layers, ANM curves exhibit a striking “mid-phase rise and end-phase collapse” pattern that matches the Ideal Refusal-Jailbreak JSS waveform projected into the same 0–1 domain. According to the results, layer-wise correlations between ANM and Ideal Refusal-Jailbreak JSS (r=0.70,0.56,0.54;RMSE=0.130.20r=0.70,0.56,0.54;\ $\text${RMSE}=0.13–0.20)[hastie2009elements] reveal a strong phase-locked coupling between representational separability and refusal intensity. In each layer, the mid-segment elevation ($s\approx 45–65$) corresponds to a surge of refusal semantics, while the late-stage decline ($s>70$) aligns with the erosion of this tendency as jailbreak success emerges.
The consistent alignment between ANM and Ideal Refusal-Jailbreak separability provides a quantitative bridge between latent-space dynamics and explicit refusal behavior, establishing an interpretable and measurable indicator of the model’s latent safety awareness.
5.2.3 Conclusion
Overall, we demonstrate that a model’s safety is manifested not only in its observable output content, but more fundamentally in the organization and evolution of its latent representation trajectories under our method. Therefore, analyzing these trajectory’s separability can equally reveal the core safety mechanisms that traditional output-based evaluations seek to capture.
5.3 Model Differentiation and Safety Characterization
This section presents the core capability of our evaluation framework: its ability to precisely and transferably characterize safety differences across models. In other words, the JSS-based system not only captures the internal representational dynamics of a model but also enables multi-scale and fine-grained quantification of safety variation and underlying regularities across different alignment levels. Within this overarching framework, our analyses proceed along two complementary axes:
  • Inter-Model Dimension (Model-Level Differentiation): examining subtle variations in latent-space separability among instruction-aligned models to reveal the continuous distribution of intermediate safety regimes;
  • Intra-Model Dimension (Weight-Level Evolution): contrasting forward DPO and inverse DPO settings to observe how safety alignment evolves directionally with changes in alignment weight.
Together, these two perspectives form a progressive precision hierarchy: from inter-model to intra-model analysis—demonstrating that the JSS framework consistently reflects alignment strength across granularities, achieving a unified, continuous, and model-invariant scale for safety evaluation.
Figure 5: Layered subset rank consistency with statistically significant (bottom right) agreement across all $\tau$ thresholds.
5.3.1 Cross-Model Validation of JSS-Derived Safety Proxies
For model-level differentiation, we collected a diverse set of models and evaluated their performance under a series of standardized red-teaming safety tests (21 testing protocols with over 7000 test cases). For each model, the results were aggregated into a single averaged safety score (Hybrid Score). Our goal in this experiment is twofold: (1) to derive proxy safety indicators from our JSS-based evaluation framework, and (2) to assess the consistency between the proxy-based ranking and the ranking obtained from conventional red-teaming safety evaluations.
Building upon the JSS evaluation indices defined earlier, we construct two higher-level proxy metrics that capture complementary aspects of safety separability.
$\mathrm{JB/PB\ Ratio}=\frac{\sum_{s\in[0,1]}\ \mathrm{JSS}_{s}\big(\mathrm{J},\mathrm{B}\big)}{\sum_{s\in[0,1]}\ \mathrm{JSS}_{s}\big(\mathrm{P},\mathrm{B}\big)}$ (27)
A higher value indicates that the overall separation strength between Jailbreak and Baseline trajectories is significant based on the metric that between PlainQuery and Baseline. This reflects stronger discriminability of jailbreak trajectories.
$\mathrm{JR\ Min/Max}=\frac{\min_{s\in[0,1]}\ \mathrm{JSS}_{s}\big(\mathrm{J},\mathrm{R}\big)}{\max_{s\in[0,1]}\ \mathrm{JSS}_{s}\big(\mathrm{J},\mathrm{R}\big)}$ (28)
A higher ratio indicates that the Ideal Refusal-Jialbreak separability remains stable across the entire trajectory. If the separation persists throughout decoding, the minimum approaches the maximum, resulting in a higher ratio. Conversely, if the separation collapses near the end—implying successful jailbreak intrusion and erosion of refusal—this ratio drops sharply, capturing the characteristic “mid-phase emergence–end-phase collapse” dynamics. To obtain a unified scalar JSS-proxy for safety alignment evaluation, we combine the two secondary metrics with equal weighting:
$\mathrm{JSS\text{-}Proxy}=0.5*\mathrm{JB/PB\ Ratio}+0.5*\mathrm{JR\ Min/Max}.$ (29)
We quantify the agreement between our proxy-based safety ranking $r_{\text{jss-proxy}}$ and the ranking obtained from conventional red-teaming evaluations $r_{\text{Hybrid}}$ using Kendall’s $\tau$ correlation coefficient:
$\tau(r_{\text{jss-proxy}},r_{\text{Hybrid}})\in[-1,1].$ (30)
Figure  5 and  3 show that cross a sufficiently large model set, the proxy-based ranking exhibits strong concordance with the red-teaming ground truth. This demonstrates that our JSS-derived proxy indicators can accurately reproduce traditional safety orderings within representative alignment regimes, while requiring far lower evaluation cost. As the model pool expands, $\tau$ shows a mild decrease but remains consistently high, with $p$-values far below conventional significance thresholds. This indicates that even under broader model coverage, the alignment between the proxy-based and red-teaming rankings remains statistically significant. Such behavior illustrates the robust hierarchical consistency of our proxy system, maintaining statistically significant agreement at scale while achieving near-perfect correspondence within focused model subsets.
(a) Inverse DPO Baseline-Jailbreak
(b) Inverse DPO Jailbreak-Plian refusual
(c) Inverse DPO Jialbreak-Refuse existing
(d) Forward DPO Baseline-Jailbreak
(e) Forward DPO Jailbreak-Plian refusual
(f) Forward DPO Jialbreak-Refuse existing
Figure 6: Evolution of JSS and behavioral safety metrics under Forward vs. Inverse DPO. The dynamic evolution of both surface-level safety behavior metrics (solid lines: unsafe rate and refusal rate) and latent proxy metrics (dashed lines: JSS ) as functions of the training global step during the Forward DPO (lower row) and Inverse DPO (upper row) processes. The three columns respectively correspond to separability relations across different dialogue families.
5.3.2 Consistency under Model Weight Variations
In essence, we demonstrate that the geometric safety scale captured by JSS not only reflects static differences among models but also dynamically responds to behavioral transitions induced by weight adjustments. Specifically, we introduce two opposite preference directions: Forward DPO and Inverse DPO, which respectively correspond to the enhancement and attenuation of safety alignment. We employ the preference optimization loss to fine-tune the same base model in two opposite directions:
$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\,\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\Big[\log\sigma\!\big(\beta\,\Delta_{\theta}\big)\Big]$ (31)
$\text{where }\;\Delta_{\theta}=\log p_{\theta}(y^{+}\!\mid\!x)-\log p_{\theta}(y^{-}\!\mid\!x).$ Specifically, we deploy it for two purposes:
Forward DPO. The model is fine-tuned with safety-oriented preferences, where compliant responses are favored over harmful ones ($y^{+}$ = safe, $y^{-}$ = harmful).
Inverse DPO. We reverse the preference direction so that harmful responses are favored over compliant ones ($y^{+}$ = harmful, $y^{-}$ = safe).
To avoid conclusions biased by a single fine-tuning intensity, we systematically generate a structured set of model variants within the LoRA parameter space, fixing the base model and dataset while varying: LoRA target modules and LoRA Rank.
Figure  6 illustrates that from a correlation perspective, the local peaks and troughs of the JSS curves strongly align with fluctuations in the surface safety metrics, including both Unsafe Rate and Refusal Rate (Definition in Appendix Section B). This observation indicates that the degree of separability in latent space evolves in synchrony with the model’s explicit refusal behaviors in output space.
A closer examination reveals a consistent time lag of several training windows between the changes in JSS and the corresponding shifts in surface metrics, with the effect being particularly evident in the early stages of Inverse DPO. Quantitatively, the lag is strongest under Inverse DPO (about 76% of runs exhibit lag). For Forward DPO, in the RefEx–JB family, 74% of trajectories show “JSS leads, refusal lags,” with a median shift of –4.4 steps, whereas in JB–Baseline the effect weakens but persists (64%, median –1.4). For the unsafe rate, both families show near-synchronous behavior (RefEx–JB: 45% lead / 48% lag; JB–Baseline centered around zero with median –0.6).
Together, these results show that declines in latent-space separability reliably precede the corresponding drops in observable safety behaviors—typically by several training windows—establishing JSS as a robust early-warning signal for impending safety degradation during training.
(a) Inverse DPO Pearson
(b) Forward DPO Pearson
(c) Inverse DPO Spearman
(d) Forward DPO Spearmsn
Figure 7: Safety Metrics (UR: Unsafe Rate; $R$: Refusual Rate) and JSS Correlation Heatmap
To complement the dual-line analysis, we compute Pearson and Spearman correlations between latent JSS signals and surface safety metrics across lower, middle, and upper layers (Figure 7). As shown in the heatmaps, the correlation signs are fully consistent across all layer groups and both alignment directions. This confirms that JSS reliably tracks the direction of safety strengthening or weakening, providing a stable latent–surface correspondence.
Figure 8: Comparative cost breakdown of traditional three-model red-teaming versus our JSS-only pipeline. For increasing red-teaming depths (T=1,5,10T={1,5,10}), the three panels report (left) total generated tokens, (middle) the relative share produced by each role—attacker ($M_{r}$), subject model ($M_{s}$), evaluator ($M_{e}$)—and (right) the per-token runtime cost.
5.4 Baseline Comparison: Cost and Efficiency of Traditional Red-Teaming vs. Our JSS-Only Pipeline
The goal of this experiment is to quantitatively compare resource consumption of two assessment pipelines under identical task scale and sample size conditions:
  • Ours. Representation-based evaluation using offline attack trajectories. We execute a single forward pass on the subject model $M_{s}$ to extract hidden states; no online generation or additional evaluator inference is required.
  • Baseline (traditional online red-team + llm-based evaluator). We run online red-teaming and safety evaluation using a PyRIT + Ollama setup, involving three model roles:
    • $M_{r}$: red-team attack generator,
    • $M_{s}$: subject model under test,
    • $M_{e}$: evaluator / safety discriminator.
Denote the total number of attack targets by $N$, the average number of turns per session by $\bar{T}$, and the average tokens per step by $\bar{L}$. We formalize token and time costs for both pipelines as follows.
For our method, only a forward pass per turn of offline attack dataset on $M_{s}$ is required and hidden states are cached, then we can get the token cost and runtime cost:
$C_{\text{token}}^{\text{Ours}}(N)\;=\;N\cdot\bar{T}_{\text{offine}},$ (32)
$C_{\text{time}}^{\text{Ours}}\;=\;\mathrm{Time}_{M_{s}^{\text{forward}}}\;+\;\mathrm{Time}_{\text{offline\_JSS}}\;+\;\mathrm{Time}_{\text{benign\_con}},$ (33)
where $\mathrm{Time}_{M_{s}^{\text{forward}}}$ denotes the cost of $C_{\text{token}}^{\text{ours}}(N,J)$, $\mathrm{Time}_{\text{offline\_JSS}}$ denotes the downstream cost for JSS statistics, and $\mathrm{Time}_{\text{benign\_con}}$ denotes time to construct the benign distribution (both $\mathrm{Time}_{\text{benign\_con}}$ and $\mathrm{Time}_{\text{offline\_JSS}}$ are typically much smaller than generation time, so we ignore them in following analysis).
For traditional baseline, each red-team attempt involves a three-way inference flow for $T$ attack turns and $N$ attack targets:
$C_{\text{token}}^{\text{Base}}(N)\;=\;N\cdot\bar{T_{r}}\cdot\big(\bar{L}_{r}\;+\;\bar{L}_{s}\;+\;\bar{L}_{e}\big),$ (34)
$C_{\text{time}}^{\text{Base}}\;=\;\mathrm{Time}_{M_{r}}\;+\;\mathrm{Time}_{M_{s}}\;+\;\mathrm{Time}_{M_{e}},$ (35)
where $\bar{L}_{r},\bar{L}_{s},\bar{L}_{e}$ denote the average token lengths for red-team generation, subject-model generation, and evaluator judgment respectively. $T_{r}$ denotes the average turn number of attacks generated by redteaming models.
We run evaluations at multiple red-team intensities $N=\{n_{1},n_{2},\dots\}$ using PyRIT’s redteaming module to generate online attacks, and record produced token counts and elapsed times. For the JSS method, we count a single forward pass per example as one token-equivalent for bookkeeping; the JSS evaluation uses the same original instruction set $N$ corresponding offline multi-turn jailbreak datasets, derived RefEx and PlainR sets. We record token-equivalent costs and elapsed wall-clock time for both pipelines.
We compare the two pipelines along the following axes:
  • Relative share of output tokens: the fraction of total system tokens generated by the subject model $M_{s}$ during evaluation.
  • Total generated-token count: the overall number of tokens produced by all system components during evaluation.
  • Per-token time cost: wall-clock runtime divided by total produced tokens; this metric measures system throughput. Lower per-token time denotes higher efficiency.
Figure 8 summarizes the cost comparison across all metrics, all roles, and all red-teaming depths, the JSS-only pipeline remains dramatically—often orders of magnitude—more efficient than the traditional three-model setup.
Together, these three findings establish that N-GLARE reduces evaluation cost by at least two orders of magnitude while preserving ranking fidelity, effectively removing the generative bottleneck that makes traditional red-teaming expensive, slow, and difficult to scale.
(a) Ranking stability across four jailbreak datasets.
(b) Ranking stability across four benign datasets.
Figure 9: Stability of normalized JSS-induced ranking value under varying sample sizes across jailbreak and benign datasets. Across all settings, the curves remain smooth and largely monotonic, confirming that N-GLARE yields stable safety ordering even when trajectory sampling budgets are perturbed.
(a) Sensitivity analysis on jailbreak datasets.
(b) Sensitivity analysis on benign datasets.
Figure 10: Sensitivity heatmaps for JSS-based ranking across datasets, models (see Appendix Table III for alphabetical labels mapping), baselines, and sample sizes. Each heatmap visualizes pairwise rank consistency between two evaluation settings computed with two rank correlation metrics (S = Spearman, K = Kendall), where darker cells indicate higher agreement. Rows and columns correspond to jailbreak datasets, models, baseline datasets, or sampling budgets $n$.
5.5 Model Sensitivity and Systemic Stability Experiments
If our method is indeed capturing an intrinsic property of the model’s alignment dynamics—rather than dataset-specific artifacts—then its induced safety ordering should remain stable when we perturb the underlying data and sampling conditions. The goal of this section is therefore to test the systemic stability of our JSS-based safety scale. Concretely, we ask:
  • whether the relative JSS differences between models (i.e., the safety ranking induced by $\mathrm{JSS}$) remain consistent when we vary which jailbreak trajectories are used under different attack datasets and how many sample size $n$ trajectories are included; and
  • whether the same stability holds when we perturb the benign reference distribution, by changing the source corpus and the number of benign samples used to estimate the natural manifold.
To assess whether JSS captures a dataset-agnostic and sample-size–agnostic property of model alignment dynamics, we evaluate stability under both jailbreak and benign perturbations. For jailbreak analysis, we use four multi-turn attack datasets—(D)=tom-gibbs, (A)=XGuard-Train, (B)=SafeMTData, (C)=MTSARed—and vary the number of included targets to produce small/medium/large (n)-sample conditions (each yielding ($5\text{–}10\times n$) activation trajectories). Figure  10 shows that across all settings, JSS-induced rankings remain highly consistent: Dataset–Dataset correlations reach mean Spearman (0.7150) and Kendall (0.5411); Model–Model comparisons show stable ordering; and varying (n) preserves strong agreement (mean Spearman (0.7973), Kendall (0.6328)). Even in pairwise ($n\times n$) evaluations, consistency remains extremely high (Spearman (0.9691), Kendall (0.8677)). An identical analysis on benign-distribution estimation—using four instruction-following corpora (LM)=LMSYS, (OA)=OASST1, (OP)=OpenHermes, (UL)=UltraChat—shows similarly robust behavior, with Dataset–Dataset correlations (mean Spearman (0.7012), Kendall (0.5237)) and Model–Model comparisons both remaining stable, and sample-size variation maintaining strong consistency (mean Spearman (0.81), Kendall (0.61)). Finally, Figure 9 provides a more intuitive visualization: as (n) increases, normalized JSS ranking curves across all jailbreak and benign datasets evolve smoothly and converge without reversals, reinforcing that JSS-based safety ordering is structurally stable across datasets, models, and sampling budgets.
6 Conclusion
This paper introduces N-GLARE, a non-generative, representation-based framework for LLM safety evaluation, showing that a model’s safety state is encoded not only in its outputs but in the geometry and dynamics of its latent trajectories. By constructing a benign distribution manifold and using JSS to quantify how jailbreak and ideal-refusal trajectories deviate from it, we obtain safety indicators that are comparable across models. Across more than 40 models and multiple red-teaming strategies, N-GLARE’s safety rankings remain consistent with large-scale red-teaming evaluations while reducing evaluation cost by over two orders of magnitude. Beyond reproducing traditional benchmarks, N-GLARE uncovers finer-grained safety dynamics, such as the phase structure of jailbreak progression, the synchronized evolution of refusal tendency, and monotonic trends under forward and inverse DPO, showcasing latent-trajectory analysis as a scalable paradigm for model-level safety diagnostics and providing a new foundation for continuous alignment monitoring and future alignment research.
Appendix A Model Safety Evaluation Results
Table LABEL:tab:jb_safety_multi presents the performance of 36 large language models tested across various security evaluation tasks. These tasks primarily measure their ability to resist jailbreak attacks, bypass protective mechanisms, and handle harmful instructions. The data in the table includes the models’ effectiveness in rejecting malicious requests, their ability to defend against bypass attacks, and their vulnerabilities in specific tasks such as ASCII art interpretation exploits and role-playing attacks.
TABLE I: Model Safety Benchmark Results
[…]
1 Llama-3.2-3B-Instruct 0.6733 0.2899 81.89 91.18 83.62 44.42 47.31 97.85 100 99.57 94.40 96.12 94.05 91.38 95.72 90.95 97.71 98.28 78.83 97.41 100 95.69 64.08 98.28
2 Qwen/Qwen2.5-7B-Instruct 0.5346 0.4414 55.17 97.78 17.42 13.82 50.28 56.94 28.93 95.26 42.81 96.88 8.29 33.69 86.80 33.16 90.24 98.63 95.09 86.51 99.52 93.10 33.91 84.74
3 mistralai/Mistral-7B-Instruct-v0.3 0.3582 0.1776 34.69 61.03 8.12 6.03 25.53 47.01 3.45 20.26 25.43 23.71 2.05 31.19 37.87 23.62 70.69 71.55 63.12 79.05 84.05 80.60 23.41 78.02
4 Llama-3.1-8B-it 0.4312 0.2538 93.22 100 99.14 65.38 100 98.71 99.43 99.14 97.63 100 97.43 89.66 99.57 95.30 98.28 100 93.63 97.41 98.71 100 90.52 100
5 google/gemma-7b-it 0.6992 0.5768 64.02 94.50 56.03 36.85 48.08 69.83 100 72.41 38.36 97.86 19.86 71.92 90.95 64.39 96.12 90.95 69.26 80.60 97.84 96.12 29.70 90.09
6 Mistral-7B-Instruct-v0.2 0.3400 0.3526 36.92 40.20 7.66 4.77 54.74 42.49 6.03 12.07 19.83 24.57 1.49 40.73 29.13 32.61 63.52 78.02 60.25 65.95 97.41 79.74 32.62 67.38
7 nvidia/Nemotron-Mini-4B-Instruct 0.5784 0.3118 51.57 96.85 22.83 12.54 95.28 50.00 99.21 29.13 35.43 91.49 7.31 38.12 93.97 33.00 85.83 99.21 92.13 82.68 96.06 74.02 23.08 75.00
8 Baichuan2-7B-Chat 0.2792 0.5432 33.17 46.17 35.04 11.50 13.79 35.97 20.16 13.96 10.84 37.72 10.03 11.74 23.35 45.94 71.80 47.86 58.18 74.25 84.05 72.33 24.07 62.12
9 gemma-2-9b-it 0.3329 0.4727 67.63 95.62 68.10 2.15 89.22 78.02 25.00 25.43 64.67 53.45 31.21 93.10 96.99 24.35 85.34 100 100 89.66 100 99.57 78.02 92.27
10 TinySwallow-1.5B-Instruct 0.6708 0.3339 68.22 98.43 44.88 14.76 98.43 71.65 100 94.49 44.88 100 46.13 63.24 98.43 57.43 96.06 98.43 97.64 86.61 100 98.43 57.14 92.13
11 microsoft/Phi-3-mini-4k-instruct 0.2995 0.5432 63.90 96.05 78.74 21.09 92.91 74.80 97.64 99.21 37.80 99.21 11.49 33.89 96.85 61.72 87.40 100 68.04 88.19 100 97.64 57.98 88.19
12 google/gemma-3-12b-it 0.6629 0.3645 71.16 98.45 85.83 18.11 85.04 90.55 21.71 88.28 88.19 86.99 10.51 64.66 96.85 63.79 88.98 97.64 100 95.28 100 96.85 67.50 96.85
13 Llama-3.1-Tulu-3-8B 0.9034 0.2734 65.95 96.26 32.81 43.19 49.68 62.89 96.25 100 38.58 98.43 10.79 56.41 88.75 73.97 100 100 90.55 100 100 100 65.54 100
14 granite-3.0-8b-instruct 0.4594 0.5180 61.30 99.63 27.59 58.79 28.63 63.39 55.17 89.37 15.19 91.81 12.89 36.07 98.71 60.83 99.14 99.14 78.15 91.81 99.14 98.71 61.55 86.21
15 DeepSeek-R1-Distill-Qwen-7B 0.5685 1.0285 53.49 65.73 25.20 28.38 82.67 75.78 26.77 83.43 64.84 83.12 33.90 49.57 44.09 89.06 59.68 54.33 58.24 87.58 33.49 77.17 42.23 87.64
16 microsoft/Phi-3-mini-128k-instruct 0.2784 0.4763 53.82 94.87 30.60 6.78 84.48 71.98 80.17 95.69 21.98 97.41 5.38 26.61 93.58 47.65 80.60 100 67.42 84.48 100 94.86 44.63 80.76
17 Mistral-NeMo-Minitron-8B-Instruct 0.7369 0.3435 47.88 84.85 19.69 21.76 35.55 48.41 14.61 43.51 29.69 36.32 3.75 43.34 91.34 36.12 90.55 86.87 74.72 87.49 96.06 81.38 59.08 71.79
18 01-ai/Yi-1.5-9B-Chat 0.1992 0.7507 39.36 85.17 12.93 3.39 39.24 61.64 27.35 9.83 21.98 53.06 11.75 23.60 67.95 27.14 74.68 86.18 47.63 77.34 93.10 75.54 27.04 76.82
19 Yi-6B-Chat 0.5100 0.3796 43.36 46.25 9.32 8.78 36.01 42.34 60.51 73.18 13.26 80.69 4.87 69.47 74.88 31.84 81.33 94.40 61.35 72.10 97.84 92.24 23.35 67.72
20 Qwen/Qwen2.5-0.5B-Instruct 0.5723 0.1777 44.05 75.06 11.71 4.45 59.99 48.15 93.13 47.45 23.24 92.67 12.81 32.20 61.16 38.57 80.76 67.81 86.92 76.76 98.28 61.45 23.84 66.08
21 Qwen2.5-1.5B-Instruct 0.6726 0.3052 52.40 98.90 11.82 17.69 85.34 48.29 99.14 72.75 29.23 97.43 14.29 23.51 87.17 43.02 95.69 95.69 91.10 81.03 99.14 85.78 21.97 76.06
22 Qwen2.5-14B-Instruct 0.4591 0.4041 68.61 99.63 28.80 4.27 91.38 83.19 78.45 98.28 58.80 98.28 36.62 58.19 93.60 67.38 96.55 100 100 94.83 96.12 99.14 77.59 92.67
23 google/gemma-3-4b-it 0.4624 0.2563 57.75 93.87 50.39 11.81 56.69 75.59 17.97 42.18 69.29 39.37 22.56 49.14 94.53 52.13 66.93 97.64 98.43 91.34 100 98.43 51.26 95.28
24 google/gemma-3-1b-it 0.6095 0.5359 65.19 96.15 59.06 15.50 60.77 73.23 83.46 22.05 70.08 80.60 44.69 41.15 96.85 37.07 71.65 100 96.85 93.70 99.21 94.49 65.55 88.98
25 gemma-2-2b-it 0.2755 0.2885 65.56 72.79 46.12 48.08 85.38 81.03 38.57 23.85 70.93 66.38 32.13 76.55 98.71 19.31 82.33 100 99.57 90.09 66.67 99.57 83.33 90.09
26 DeepSeek-R1-Distill-Qwen-1.5B 0.2589 0.7154 40.60 41.18 19.53 30.10 61.60 82.36 27.09 31.92 66.89 15.25 17.74 49.94 25.39 38.26 39.14 25.30 44.60 63.69 34.37 49.74 51.52 60.61
27 DeepSeek-R1-Distill-Qwen-14B 0.5352 1.4011 57.53 76.38 34.65 24.22 80.60 75.59 32.81 27.34 70.08 25.20 34.78 65.52 59.06 74.05 74.21 44.09 80.31 84.72 77.07 88.98 63.03 83.20
28 deepseek-ai/DeepSeek-R1-Distill-Llama-8B 0.5352 1.4011 51.17 70.25 21.26 23.26 92.23 61.71 39.84 36.60 49.22 31.25 29.42 53.45 51.97 86.11 69.29 59.84 62.20 86.03 55.35 85.04 39.50 79.42
29 meta-llama/Meta-Llama-3-8B-Instruct 0.0798 0.2441 79.98 91.61 71.12 80.26 72.13 97.84 100 71.37 93.99 100 20.84 83.33 96.57 97.13 93.16 67.24 98.71 98.71 99.57 100 87.60 98.71
30 meta-llama/Llama-2-7b-chat-hf 0.7188 0.4138 64.47 46.79 98.71 7.73 99.14 88.53 100 89.80 47.09 98.71 12.33 73.07 97.41 58.40 91.38 100 76.60 90.09 100 99.14 85.99 88.36
31 meta-llama/Llama-2-13b-chat-hf 0.7386 1.4244 75.40 63.94 98.56 85.46 83.73 80.52 99.71 98.28 59.48 99.43 12.36 60.51 96.57 66.72 99.14 99.57 75.49 91.38 99.57 100 89.66 90.99
32 Qwen/Qwen2.5-72B-Instruct 0.5530 0.6269 79.41 97.71 54.61 10.80 98.86 78.06 100 98.58 88.67 95.74 83.13 61.76 98.29 86.25 95.73 100 98.29 87.75 99.43 98.29 81.63 92.88
33 Qwen/Qwen2.5-32B-Instruct 0.4748 0.5229 68.81 97.10 56.56 12.57 94.17 79.12 44.30 92.87 55.67 97.69 28.12 64.47 96.24 64.71 93.44 99.60 97.81 92.24 100 97.48 70.77 89.43
34 Qwen/Qwen2.5-3B-Instruct 0.3775 0.3692 51.16 98.92 23.10 9.36 36.22 49.57 63.68 81.47 25.64 97.41 11.10 26.18 94.45 28.74 93.10 99.14 91.02 78.11 100 95.26 35.22 80.60
35 meta-llama/Meta-Llama-3-70B-Instruct 0.9358 0.2999 59.43 74.28 93.10 40.14 14.35 92.75 100 96.84 17.28 83.49 51.15 45.19 91.52 83.43 81.52 94.89 68.37 57.77 54.30 94.01 45.48 55.94
36 Llama-3.1-Nemotron-70B-Instruct-HF 0.2293 0.4283 69.69 92.96 32.28 84.15 79.31 75.78 84.37 66.93 54.33 78.74 14.30 60.34 94.83 75.86 95.28 96.06 89.06 94.53 100 98.43 69.23 93.16
TABLE II: Model–Abbreviation Mapping of Subset models Scatter Figure
Model Abbr
01-ai/Yi-1.5-9B-Chat CHA-9B
Baichuan2-7B-Chat BAI-7B
DeepSeek-R1-Distill-Qwen-1.5B DEE-1.5B
DeepSeek-R1-Distill-Qwen-14B DEE-14B
DeepSeek-R1-Distill-Qwen-7B DEE-7B
Llama-3.1-8B-it LLA-8B
Llama-3.1-Nemotron-70B-Instruct-HF LLA-70B
Llama-3.1-Tulu-3-8B LLA-8B
Llama-3.2-3B-Instruct LLA-3B
Mistral-7B-Instruct-v0.2 MIS-7B
Mistral-NeMo-Minitron-8B-Instruct MIS-8B
Qwen/Qwen2.5-0.5B-Instruct INS-0.5B
Qwen/Qwen2.5-32B-Instruct INS-32B
Qwen/Qwen2.5-3B-Instruct INS-3B
Qwen/Qwen2.5-72B-Instruct INS-72B
Qwen/Qwen2.5-7B-Instruct INS-7B
Qwen2.5-1.5B-Instruct QWE-1.5B
Qwen2.5-14B-Instruct QWE-14B
SakanaAI/TinySwallow-1.5B-Instruct INS-1.5B
Yi-6B-Chat YIX-6B
deepseek-ai/DeepSeek-R1-Distill-Llama-8B BXX-8B
gemma-2-2b-it GEM-2B
gemma-2-9b-it GEM-9B
google/gemma-3-12b-it ITX-12B
google/gemma-3-1b-it ITX-1B
google/gemma-3-4b-it ITX-4B
google/gemma-3-7b-it ITX-7B
granite-3.0-8b-instruct GRA-8B
meta-llama/Llama-2-13b-chat-hf HFX-13B
meta-llama/Llama-2-7b-chat-hf HFX-7B
meta-llama/Meta-Llama-3-70B-Instruct INS-70B
meta-llama/Meta-Llama-3-8B-Instruct INS-8B
microsoft/Phi-3-mini-128k-instruct INS-128K
microsoft/Phi-3-mini-4k-instruct INS-4K
mistralai/Mistral-7B-Instruct-v0.3 VXX-7B
nvidia/Nemotron-Mini-4B-Instruct INS-4B
TABLE III: Model ID Mapping with Alphabetical Labels
Model ID Letter
Llama3-8B A
gemma3-4b B
gemma2-9b C
qwen3-8b D
qwen2.5-7b E
deepseek-llama8b F
mistral-7bv3 G
synlogic-7b H
deepseek-qwen1.5b I
Llama2-7B J
Appendix B Unsafe Rate and Refusual Rate Definition
Unsafe Rate. This metric is computed using the safety classifications provided by Qwen3Guard. A model output is counted as unsafe whenever the guard model labels it as “Unsafe.” Formally,
$\text{Unsafe Rate}=\frac{N_{\text{Unsafe}}}{N_{\text{Total}}},$
where $N_{\text{Unsafe}}$ denotes the number of responses flagged as Unsafe by the guard model, and $N_{\text{Total}}$ denotes the total number of model responses evaluated.
Refusal Rate. This metric is based on refusal-style keyword matching. A model response is counted as a refusal if it contains typical refusal patterns such as “I’m sorry”, “I cannot comply”, “I refuse to answer”, or similar formulations. Formally,
$\text{Refusal Rate}=\frac{N_{\text{Refusal}}}{N_{\text{Total}}},$
where $N_{\text{Refusal}}$ denotes the number of responses that match the refusal keyword patterns.
Complementarity. These two metrics capture opposite behavioral tendencies and are therefore complementary. Lower Unsafe Rate indicates stronger safety, while lower Refusal Rate indicates better helpfulness and less over-defensiveness. A high Unsafe Rate signals safety failures, whereas a high Refusal Rate reflects an overly conservative model that rejects too many user requests.

### v2

@@
 It characterizes hidden layer dynamics by analyzing the APT (Angular-Probabilistic Trajectory) of latent representations and introducing the JSS (Jensen-Shannon Separability) metric.
-Experiments on over 40 models and 20 red teaming strategies demonstrate that the JSS metric exhibits high consistency with the safety rankings derived from Red Teaming.
-N-GLARE reproduces the discriminative trends of large-scale red-teaming tests at less than 1% of the token cost and the runtime cost, providing an efficient output-free evaluation proxy for real-time diagnostics.
+Experiments on over 40 models and 20 red teaming strategies demonstrate that the JSS metric exhibits high consistency with Red Teaming safety rankings at less than 1% token and runtime cost.
+N-GLARE: An Non-Generative Latent Representation-Efficient LLM Safety Evaluator
+Zheyu Lin1, Jirui Yang2, Yukui Qiu3 Hengqi Guo2, Yubing Bao2, Yao Guan2 1University of California, Riverside, 2Fudan University, 3South China University of Technology Zheyu Lin: 1298lzy@gmail.com Jirui Yang: yangjr23@m.fudan.edu.cn Yukui Qiu: 202162311356@mail.scut.edu.cn Hengqi Guo: hqguo22@m.fudan.edu.cn Yubing Bao: ybbao23@m.fudan.edu.cn Yao Guan: yguan25@m.fudan.edu.cn
 1 Introduction
-In recent years, Large Language Models (LLMs) have developed rapidly, with their capability boundaries continuously expanding.
-They have been deployed in a wide range of application scenarios, exerting substantial impact on social production and daily life [grattafiori2024llama, zhou2024large].
-However, alongside this growth in capability, security risks and alignment issues have become increasingly prominent, emerging as a critical bottleneck for large-scale, safety-critical deployment [das2025security, dong2024attacks].
-Trained on vast and often uncurated Internet corpora, LLMs can inadvertently internalize social biases and potentially harmful knowledge.
-Without robust safety alignment, they may produce unsafe outputs when exposed to malicious prompts, complex contexts, or specific triggers—for example, generating cyberattack code [wei2023jailbroken], spreading misleading information [russinovich2025great], or guiding embodied agents toward physically harmful behaviors [yang2025concept].
-Ensuring that LLMs remain aligned with human values and safety norms is therefore a central challenge.
+Large Language Models must be aligned to ensure their behavior consistently follows human values and safety norms, rather than merely maximizing task performance Das et al.
+(2025); Dong et al.
+(2024).
+Without such alignment, models may generate harmful, biased, or unsafe outputs when facing adversarial prompts or complex contexts Wei et al.
+(2023); Russinovich et al.
+(2025); Yang et al.
+(2025).
 Evaluating model safety is a foundational step toward achieving such alignment.
-The dominant paradigm in both industry and academia is currently Red Teaming [lin2025against], which simulates adversarial user behavior by designing diverse, subtle attack prompts to probe the model and elicit unsafe outputs.
+The dominant paradigm in both industry and academia is currently Red Teaming Lin et al.
+(2025), which simulates adversarial user behavior by designing diverse, subtle attack prompts to probe the model and elicit unsafe outputs.
 While this online, behavior-based testing has proved effective for discovering vulnerabilities, it suffers from two fundamental limitations.
@@
 A refusal that looks safe at the surface level may simply be triggered by brittle pattern matching, rather than reflecting a robust internal representation of harmfulness.
-To move beyond purely behavioral stress testing, recent work has begun to analyze LLMs in their latent representation space [zou2023representation, li2025revisiting].
-Representation engineering shows that when a model processes different concepts (including abstract notions such as “truthfulness,” “refusal,” or “harmfulness”), its hidden activations form structured, often linearly separable patterns.
-Let $h_{l,p}\in\mathbb{R}^{d}$ denote the activation at layer $l$ and position $p$.
-For two contrasting semantic sets $\mathcal{S}^{+}$ and $\mathcal{S}^{-}$, their mean activations differ along concept directions of the form
-$o\approx\mathbb{E}[H(\mathcal{S}^{+})]-\mathbb{E}[H(\mathcal{S}^{-})],$ (1)
-suggesting that high-level semantics are explicitly encoded in latent geometry.
-Existing work primarily uses such directions at the prompt level, for example to steer individual generations or to classify the risk of a specific input.
+To move beyond purely behavioral stress testing, recent work analyzes LLMs in their latent representation space Zou et al.
+(2023); Li et al.
+(2025b).
+Representation engineering shows that different semantic concepts—such as truthfulness, refusal, or harmfulness—correspond to structured patterns in hidden activations, and contrasting concept sets can be distinguished by consistent directions in this space, suggesting that high-level semantics are explicitly encoded in latent geometry.
+Existing representation analysis work primarily uses such directions at the prompt level, for example to steer individual generations or to classify the risk of a specific input.
 However, these techniques are typically model-specific and do not directly provide a model-level safety indicator that is comparable across architectures and checkpoints.
-In this paper, we take a step in that direction and ask: can we evaluate model safety without generating full outputs, solely by examining how its internal representations evolve under different safety conditions?
-To this end, we propose N-GLARE (A Non-Generative, Latent Representation-Efficient LLM Safety Evaluator), a framework that operates entirely in latent space and bypasses token-level generation.
+In this paper, we take a step in that direction and ask: Can we evaluate model safety without generation, using internal representations that correlate with red-teaming robustness?
+To this end, we propose N-GLARE (A Non-Generative, Latent Representation-Efficient LLM Safety Evaluator).
 For each model, we construct Angular-Probabilistic Trajectories (APT) of hidden representations under different probing conditions: benign ($B$), jailbreak ($J$), plainquery ($P$) and an idealized refusal counterfactual ($R$).
@@
 We then introduce Jensen–Shannon Separability (JSS), which compares the slice-wise distributions of these turning angles across conditions, and aggregates them into model-level safety indicators, introducing other JSS-based safety scores.
-Figure 1: JSS Shows Stronger Safety Alignment in Qwen3-4B-SafeRL Compared to Base and Alignment-Removed Variants.
-Statistics on the JSS distribution for 100 jailbreak trajectories and the ideal refusal pattern.
-Figure 1 illustrates the core intuition: when comparing multiple variants of the same base model (e.g., RL-aligned, base, and alignment-removed versions of Qwen3-4B), we observe that better-aligned variants exhibit more pronounced geometric separation between benign and jailbreak trajectories in latent space, and their jailbreak trajectories are also further away from the ideal refusal pattern.
-This suggests that internal trajectory separability can serve as a proxy for the underlying safety state of the model, potentially reproducing the discriminative power of large-scale Red Teaming at a fraction of the cost.
-We formalize this idea in Section 3 by casting safety evaluation as a problem of comparing representation trajectories under $(B,J,R,P)$ conditions.
-We then instantiate our method in Section 4: benign activations define a whitened manifold, the trajectory turning angle encodes instantaneous deviation from this manifold, and Jensen–Shannon divergence over angle distributions along the progress axis yields JSS metrics that are both geometrically interpretable and amenable to cross-model comparison.
+Figure 1: JSS exhibits clear separability across safety states while achieving orders-of-magnitude higher token efficiency than red-teaming baselines.
+The core intuition is that: internal trajectory separability can serve as a proxy for the underlying safety state of the model, potentially reproducing the discriminative power of large-scale Red Teaming at a fraction of the cost.
+Figure 1 supports the core intuition: when comparing multiple variants of the same base model (e.g., RL-aligned, base, and safety-removed versions of Qwen3-4B), we observe that better-aligned variants exhibit more pronounced geometric separation between different trajectories in latent space.
 Experiments on over 40 models and 20 red-teaming strategies demonstrate that N-GLARE faithfully reproduces the safety rankings induced by traditional Red Teaming while using less than 1% of the token and runtime cost, making it suitable for agile diagnostics after training new models or updating safety mechanisms.
@@
 2.1 LLM Security and Red Teaming
-Despite the widespread implementation of safety alignment in contemporary mainstream LLMs, they continue to face extensive security challenges in practical application [Cui2024Risk].
-Attackers frequently employ meticulously designed adversarial inputs to bypass the model’s safety guardrails [zhu2023promptbench, paulus2024advprompter].
-Through so-called “jailbreak attacks,” models are induced to circumvent their established policies, consequently generating harmful content.
-Such content may be exploited for academic misconduct [oliveira2025human], facilitating cyberattacks [xu2024autoattacker], or infringing upon copyright [xu2024llms].
-Consequently, ensuring that model outputs adhere to ethical norms and effectively mitigating potential risks have become indispensable research directions in the field of artificial intelligence.
-In this context, the security evaluation of large models is paramount, wherein Red Teaming has emerged as a core methodology for assessing model safety.
-Red Teaming is an adversarial evaluation method centered on proactively discovering and identifying potential model vulnerabilities by simulating attacker strategies and behaviors.
-Existing Red Teaming approaches can be broadly categorized into two types.
-The first focuses on content safety in single-turn interactions, such as leveraging algorithms like Monte Carlo Tree Search (MCTS) and evolutionary algorithms to optimize attack prompts [sadasivan2024fast], or generating jailbreak templates through the mutation of seed prompts [yu2023gptfuzzer].
-The second category addresses security risks in multi-turn dialogue scenarios, typically employing specialized agents for attack planning, executing dynamic multi-turn jailbreaks, and evaluating the attack’s efficacy [rahman2025x, weng2025foot, ren2025llms].
-Although existing methods show a trend towards diversification, their mainstream paradigm still generally follows a three-stage process: first, constructing the attack prompt; second, obtaining the target model’s output for the prompt; and finally, judging the success of the attack based on the evaluation of the model’s output results.
-This model leads to particularly high evaluation overhead for Red Teaming.
+Red teaming evaluates LLM safety by simulating attacker strategies to proactively identify and probe model vulnerabilities.Existing Red Teaming approaches can be broadly categorized into two types.
+The first focuses on content safety in single-turn interactions, such as leveraging algorithms like Monte Carlo Tree Search (MCTS) and evolutionary algorithms to optimize attack prompts Sadasivan et al.
+(2024), or generating jailbreak templates through the mutation of seed prompts Yu et al.
+(2023).
+The second category addresses security risks in multi-turn dialogue scenarios, typically employing specialized agents for attack planning, executing dynamic multi-turn jailbreaks, and evaluating the attack’s efficacy Rahman et al.
+(2025); Weng et al.
+(2025); Ren et al.
+(2025).
+Existing methods, despite increasing diversification, largely follow a three-stage pipeline that includes attack prompt construction, target model generation, and output-based evaluation, which inherently incurs high overhead in red-teaming assessments.
 2.2 LLM Latent Space Analysis Techniques
@@
 Embedding analysis reveals semantic meaning by examining the static geometry of latent space vectors.
-Common methods like cosine similarity ($Sim(\vec{A},\vec{B})=\frac{\vec{A}\cdot\vec{B}}{\|\vec{A}\|\|\vec{B}\|}$) and dimensionality reduction visualization like t-SNE can reveal semantic clusters (e.g., “king” with “queen”) [rogers2020primer].
-Furthermore, domain-adaptation analysis shows that fine-tuning (e.g., BioBERT [lee2020biobert]) reshapes the latent space, forming more refined domain-specific subspaces to enhance downstream task performance.
+Common methods like cosine similarity ($Sim(\vec{A},\vec{B})=\frac{\vec{A}\cdot\vec{B}}{\|\vec{A}\|\|\vec{B}\|}$) and dimensionality reduction visualization like t-SNE can reveal semantic clusters (e.g., “king" with “queen") Rogers et al.
+(2020).
+Furthermore, domain-adaptation analysis shows that fine-tuning such as BioBERT Lee et al.
+(2020) reshapes the latent space, forming more refined domain-specific subspaces to enhance downstream task performance.
 2.2.2 Probing and Intervention
 This technique shifts from passive analysis to active decoding and control.
-Early probing [tenney2019bert] used simple linear classifiers ($y=f(H)$) to test if latent activations $H$ encoded specific linguistic properties, discovering that information is stored hierarchically in LLMs.
-A more advanced paradigm, Representation Engineering (RepE) [zou2023representation], controls model behavior by directly manipulating latent representations.
+Early probing Tenney et al.
+(2019) used simple linear classifiers ($y=f(H)$) to test if latent activations $H$ encoded specific linguistic properties, discovering that information is stored hierarchically in LLMs.
+A more advanced paradigm, Representation Engineering (RepE) Zou et al.
+(2023), controls model behavior by directly manipulating latent representations.
 Its core is identifying vector directions, or concept operators $o^{c}$, that represent high-level concepts (e.g., truthfulness).
-This is often achieved by contrasting the activation differences from different input sets ($\mathcal{S}^{c+}$ vs $\mathcal{S}^{c-}$):
-$o^{c}\approx\mathbb{E}[H(\mathcal{S}^{c+})]-\mathbb{E}[H(\mathcal{S}^{c-})]$ (2)
-Here, $H(\mathcal{S})$ represents the mean activation of model $\mathcal{M}$ when processing a specific input set $\mathcal{S}$.
-RepE is widely applied in LLM safety [cai2024self, wang2025model, wang2024inferaligner, cao2024nothing], for instance, by using positive or negative steering vectors to enhance truthfulness [marks2023geometry, li2023inference, wang2025adaptive, rimsky2024steering] or suppress meta-behaviors like “refusal” [arditi2024refusal].
-2.2.3 Attention Mechanism Analysis
-Attention mechanism analysis interprets the model’s information flow via its weight distribution $A=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)$.
-Here, the $Q,K,V$ matrices are linear projections of the layer input $h_{l-1}$.
-Heatmap visualization [vig2019multiscale] reveals that attention exhibits hierarchical properties (from syntax to semantics across layers).
-At the same time, research [michel2019sixteen] indicates high redundancy in attention mechanisms.
-To address its computational complexity, models like Perceiver [jaegle2021perceiver] introduce low-dimensional latent variables to compress the computation.
+This is often achieved by contrasting the activation differences from different input sets ($\mathcal{S}^{c+}$ vs $\mathcal{S}^{c-}$): $o^{c}\approx\mathbb{E}[H(\mathcal{S}^{c+})]-\mathbb{E}[H(\mathcal{S}^{c-})]$ Here, $H(\mathcal{S})$ represents the mean activation of model $\mathcal{M}$ when processing a specific input set $\mathcal{S}$.
+RepE is widely applied in LLM safety Cai et al.
+(2024); Wang et al.
+(2025a, 2024); Cao et al.
+(2024), for instance, by using positive or negative steering vectors to enhance truthfulness Marks and Tegmark (2023); Li et al.
+(2023); Wang et al.
+(2025b); Rimsky et al.
+(2024) or suppress meta-behaviors like “refusal" Arditi et al.
+(2024).
 2.2.3 Latent Space Dynamic Analysis
 Dynamic analysis focuses on the activation trajectories during the reasoning process.
-Researchers not only track attention changes during Chain-of-Thought (CoT) [wei2022chain] but also begin to analyze the latent space trajectories themselves.
+Researchers not only track attention changes during Chain-of-Thought (CoT) Wei et al.
+(2022) but also begin to analyze the latent space trajectories themselves.
 For example, Li et al.
-[li2025core] proposed Chain-of-Reasoning Embedding (CoRE), which treats the reasoning process as a series of hidden states $h_{l,p_{T}}$.
-They found that redundant reasoning (”overthinking”) manifests as identifiable cyclical fluctuations in the trajectory.
-Based on this, the CoRE-Eval framework was developed, enabling label-free self-evaluation to dynamically terminate inefficient reasoning, thereby improving efficiency.
+Li et al.
+(2025a) proposed Chain-of-Reasoning Embedding (CoRE).
+They found that redundant reasoning manifests as identifiable cyclical fluctuations in the trajectory.
+Based on this, it enables label-free self-evaluation to dynamically terminate inefficient reasoning, thereby improving efficiency.
 Despite significant progress in analysis techniques for LLM latent spaces, the field still faces the challenge of model-specificity: the identified features, concept operators (like $o^{c}$), attention heads, or reasoning trajectories are often highly dependent on a single model $\mathcal{M}$ and lack transferability across models.
-3 Problem Statement
-Safety evaluation of large language models is currently dominated by generative red-teaming: a model is queried with adversarial prompts generated by red-teaming LLMs, asked to produce multi-token outputs, and then judged by a human or a separate LLM-based evaluator.
-This approach is expensive (because evaluation cost grows with generated tokens), noisy (because harmfulness must be inferred from surface text), and incomplete (because it reveals only the final behavior, not the underlying safety state).
-Formally, let $M_{\text{subject}},M_{s}$ be a model and let $x$ be a potentially harmful input drawn from an adversarial distribution $\mathcal{D}_{\mathrm{adv}}$ or online $M_{\text{attacker}},M_{r}$.
-A generative evaluator, another LLM $M_{\text{evaluator}},M_{e}$ observes the decoded output $y=M_{s}(x)$ and produces an assessment $\mathrm{Eval_{M_{e}}}(y)$ that approximates the latent safety function $f_{\mathrm{safety}}(M_{s},x)$.
-That is,
-$\mathrm{Eval_{M_{e}}}(y)\approx f_{\mathrm{safety}}(M_{s},x),$
-where this 0/1 function, $f_{\mathrm{safety}}$ represents whether the model processes $x$ in a safe or unsafe manner.
-This evaluation pipeline is entirely output-dependent (by analyzing the natural language text) and therefore cannot access $f_{\mathrm{safety}}(M_{s},x)$ directly.
-In contrast, a non-generative evaluator also seeks to estimate $f_{\mathrm{safety}}(M_{s},x)$ without decoding any output tokens.
-Instead, it relies only on information available inside the $M_{s}$.
-The goal is to design a mapping
-$g:(M_{s},x)\;\mapsto\;f_{\text{safety}}(M_{s},x),$
-where $g(M,x)$ is a model-level safety signal that: (i) approximates the same generative-based quantity $\mathrm{Eval_{M_{e}}}(y)$ evaluated by LLM-based evaluator, (ii) is comparable across models and prompts, and (iii) does not rely on online red-teaming text generation or attackers $M_{r}$.
-The core problem addressed in this work is therefore:
-How can we construct a non-generative safety metric $g(\cdot,\cdot)$ that reliably reflects a model’s internal safety state and correlates with its behavioral robustness observed under traditional red-teaming?
-This paper develops such a metric by operating entirely in the representation space, enabling low-cost, model-level safety evaluation without generation.
 3 Methodology
-We now instantiate the representation-space safety evaluation framework outlined above.
-Given a set of models $\mathcal{M}=\{M_{1},\dots,M_{K}\}$ and probing conditions $A\in\{B,J,R,P\}$ (benign, jailbreak, ideal refusal), our goal is to map internal activation trajectories to scalar safety indicators.
-The pipeline proceeds in three stages: (i) extracting layer-wise hidden states and normalizing them along a standardized progress axis; (ii) fitting a benign manifold and defining a geometric turning angle that measures instantaneous deviation from this manifold; and (iii) aggregating slice-wise Jensen–Shannon divergences of turning-angle distributions into Jensen–Shannon Separability (JSS) metrics and a normalized safety score.
-4.1 Trajectory Extraction and Standardized Progress
-For each model $M$ and probing condition $A\in\{B,J,R,P\}$ (defined in Section 4.3), we construct probing sequences that form either genuine interaction histories (e.g., multi-turn dialogues) or canonical prefix expansions for single-turn prompts.
-At each node $t$ in such a sequence, we identify a fixed decision-making cross-section, chosen in our experiments as the point immediately before the model generates the next assistant response.
-A single forward pass at node $t$ yields the last-token hidden activations across all transformer layers,
+This section formalizes the representation-space safety evaluation framework proposed above.
+Our goal is to map internal activation trajectories to scalar safety metrics.
+We first define the notation and probing conditions, followed by the framework’s three core stages: (i) trajectory extraction and standardization; (ii) benign manifold fitting and geometric turning angle calculation; and (iii) distribution divergence measurement and safety score aggregation.
+3.1 Preliminaries: Symbols and Probing Conditions
+Given a model set $\mathcal{M}=\{M_{1},\dots,M_{K}\}$, we analyze internal representation behaviors under varying inputs.
+We define the probing condition $A\in\{B,J,R,P\}$ to correspond to four interaction history types:
+Benign ($B$): Standard task instructions conforming to natural distributions, used to fit the benign reference manifold.
+Jailbreak ($J$): Adversarial sequences designed to induce harmful outputs.
+Ideal Refusal ($R$): Identical to $J$ in user content (retaining harmful intent) but forcing the assistant response to a policy-compliant refusal template.
+This represents the model’s ideal defensive state against attacks.
+Plain Query ($P$): Direct presentation of harmful intent (without disguise) to capture the model’s unconstrained internal response to harmful information.
+To enhance cross-model comparability and reduce dimensionality, we partition Transformer layers into groups $G\in\{\text{lower},\text{middle},\text{upper}\}$ and average within each group to obtain group-level representations $h^{(A,G)}$.
+3.2 Stage 1: Trajectory Extraction and Standardization
+For a model $M$ and condition $A$, we construct probing sequences.
+At each time step $t$, we extract the hidden states of the last token across all layers at the decision-making cross-section immediately preceding the next assistant response generation:
 $\Phi_{M}^{(A)}(t)=\bigl(h_{L,t}^{(A)},\dots,h_{1,t}^{(A)}\bigr),\quad h_{\ell,t}^{(A)}\in\mathbb{R}^{d_{\ell}},$ (1)
-where $L$ is the number of layers and $d_{\ell}$ is the hidden dimension at layer $\ell$.
-To improve cross-model comparability and reduce dimensionality, we partition layers into a small number of layer groups $G\in\{\text{lower},\text{middle},\text{upper}\}$ and average within each group, obtaining group-level representations
-$h_{t}^{(A,G)}\in\mathbb{R}^{d},$ (4)
-with a common dimension $d$. For each condition $A$ and group $G$, the sequence $\{h_{t}^{(A,G)}\}_{t=1}^{T_{A}}$ defines a trajectory in latent space.
-The length $T_{A}$ of a probing sequence may vary across prompts and models.
-To enable comparison on a common timeline, we normalize each trajectory by its arc length and define a standardized progress variable $s_{t}\in[0,1]$:
-$s_{1}=0,\qquad s_{t}=\frac{\sum_{k=1}^{t-1}\|h_{k+1}^{(A,G)}-h_{k}^{(A,G)}\|_{2}}{\sum_{k=1}^{T_{A}-1}\|h_{k+1}^{(A,G)}-h_{k}^{(A,G)}\|_{2}},\quad t>1.$ (5)
-This parametrization treats trajectories from different prompts and models as curves evolving along the shared progress axis $s\in[0,1]$.
-In single-turn settings where there is no natural temporal dimension, we construct pseudo-sequences via canonical prefix augmentation (e.g., from an empty system prompt to the full user query) and apply the same normalization.
-We denote the resulting standardized trajectory under condition $A$ and group $G$ by
-$\Gamma_{M}^{(A,G)}=\{(h_{t}^{(A,G)},s_{t})\}_{t=1}^{T_{A}}.$ (6)
-4.2 Benign Manifold and Geometric Turning Angle
-To quantify deviations from normal behavior, we first construct a benign manifold using representations collected under the benign condition $B$.
+where $L$ is the number of layers and $d_{\ell}$ is the hidden dimension.
+As defined, these are aggregated into group-level representations $h_{t}^{(A,G)}\in\mathbb{R}^{d}$.
+Thus, for condition $A$ and group $G$, the sequence $\{h_{t}^{(A,G)}\}_{t=1}^{T_{A}}$ defines a trajectory in latent space.
+Since generation lengths $T_{A}$ vary across prompts and models, we define an arc-length standardized progress variable $s_{t}\in[0,1]$ to align trajectories on a unified timeline.
+For $t>1$ (with $s_{1}=0$):
+$s_{t}=\frac{\sum_{k=1}^{t-1}\|h_{k+1}^{(A,G)}-h_{k}^{(A,G)}\|_{2}}{\sum_{k=1}^{T_{A}-1}\|h_{k+1}^{(A,G)}-h_{k}^{(A,G)}\|_{2}}$ (2)
+This parameterization treats trajectories as curves evolving along a shared progress axis $s$. The standardized trajectory is denoted as:
+$\Gamma_{M}^{(A,G)}={(h_{t}^{(A,G)},s_{t})}_{t=1}^{T_{A}}.$ (3)
+Figure 2: Illustration of latent representation trajectories relative to the benign manifold.
+Solid curves denote a normally aligned model, while dashed curves indicate a jailbroken model.
+Red represents explicit refusal, yellow/orange denotes jailbreak behavior, and green corresponds to benign dialogue.
+3.3 Stage 2: Benign Manifold and Geometric Turning Angle
+To quantify deviations from normal behavior during inference, we introduce geometric metrics.
+First, we construct a benign manifold using representations collected under benign conditions ($B$) as a reference for normal generation (construction details in Appendix C).
+For any given trajectory, we examine two local directions at each step:
+Trajectory Tangent ($\tau_{t}^{(A,G)}$): Reflects the direction of representation evolution during inference.
+Outward Normal ($n^{(G)}$): Indicates the direction of maximal deviation from the benign manifold.
+We define the geometric turning angle as the angle between these vectors:
+$\theta_{t}^{(A,G)}=\arccos\left(\frac{\tau_{t}^{(A,G)\top}n^{(G)}\bigl(h_{t}^{(A,G)}\bigr)}{\|\tau_{t}^{(A,G)}\|_{2},\|n^{(G)}\bigl(h_{t}^{(A,G)}\bigr)\|_{2}}\right).$ (4)
+Intuitively, $\theta_{t}^{(A,G)}$ measures whether the trajectory trends toward or away from the benign region.
+Smaller angles indicate evolution away from the benign manifold, signaling potential safety risk shifts.
+Figure 3: APT trajectories collected from lower/middle/upper(up to down) layer groups.
+left column correspond to safety-aligned models, while right column ones are from jailbreak-tuned models.
+Safety-aligned LLMs exhibit well-separated and low-variance APT trajectories, whereas jailbreak-tuned LLMs show weakened separability and substantially higher variance across all query types, indicating degraded safety boundaries.
+Fig.
+2 further provides an intuitive three-dimensional geometric view of how these trajectories bend and diverge relative to the benign manifold.
+3.4 Stage 3: Slicing and JSS Metrics Calculation
+We quantify safety by analyzing the divergence of turning angle distributions across conditions.
+We discretize the standardized progress axis $s\in[0,1]$ into $I$ slices $\{s_{i}\}_{i=1}^{I}$.
+For each slice $s_{i}$, we collect the corresponding angles:
+$\Theta^{(A,G)}(s_{i})={\theta_{t}^{(A,G)}\mid s_{t}\in\text{slice }i}.$ (5)
+Within each slice $s_{i}$ and group $G$, we estimate the empirical distributions for two conditions (e.g., $J$ vs. $B$, or $J$ vs. $R$) and compute the Jensen–Shannon divergence (JSD):
+$\mathrm{JS}_{i}^{(G)}(A,B)=\mathrm{JS}\bigl(\mathrm{D}(\Theta_{A}^{(G)}(s_{i})),\big|,\mathrm{D}(\Theta_{B}^{(G)}(s_{i}))\bigr).$ (6)
+To obtain a model-level separability measure, we aggregate JSDs across all slices and layer groups to define the Jensen-Shannon Separability (JSS):
+$\mathrm{JSS}(A,B)=\frac{1}{|G|}\sum_{G}\left(\frac{1}{I}\sum_{i=1}^{I}\mathrm{JS}_{i}^{(G)}(A,B)\right).$ (7)
+Algorithm 1 summarizes the core computation for a single model.
+3.5 Insights: Trajectory Separability as a Latent Safety Signal
+Separability as an Indicator of Safety Awareness and Refusal Tendency. From this geometric perspective, two JSS relations hold key safety implications:
+$\mathrm{JSS}(J,B)$ (Jailbreak vs.
+Benign): Measures whether the model internally distinguishes adversarial inputs from normal dialogue.
+High separability implies acute latent risk recognition (representation shift details in Appendix B.2).
+$\mathrm{JSS}(J,R)$ (Jailbreak vs.
+Ideal Refusal): Reflects the model’s intrinsic refusal tendency and resistance to late-stage safety erosion.
+We observe a significant correlation between this metric and the refusal tendency score (ANM) introduced in Appendix B.3.
+Necessity of Trajectory-Level Aggregation.
+These observations suggest that endpoint or single-step metrics cannot jointly capture both early risk recognition and late refusal stability.
+Thus, JSS serves as a compact scalar summarizing the latent safety state across the entire inference trajectory, enabling safety evaluation without reliance on text generation.
+4 Experiments and Analysis
+In this section, we employ quantified safety feature metrics to evaluate various aligned models and compare their rankings’ correlation and consistency with those obtained from systematic traditional red-teaming evaluations, demonstrating a high degree of consistency in safety scales as well as a significant advantage in evaluation cost.
+Finally, we investigate the statistical consistency of the ranking results across different datasets and sample sizes, analyzing the stability of our evaluation framework under changes in data distribution and sampling conditions.
+Figure 4: Multi-perspective cross-model safety ranking consistency between N-GLARE and traditional red-teaming.
+Detailed figures and datapoints at Appendix Fig.
+10 and Appendix Table LABEL:tab:jb_safety_benchmark_split_A  LABEL:tab:jb_safety_benchmark_split_B.
+4.1 Model Differentiation and Safety Characterization
+This section presents the core capability of our evaluation framework: its ability to precisely and transferably characterize safety differences across models.
+Within this overarching framework, our analyses proceed along two complementary axes: Inter-Model Dimension (Model-Level Differentiation): examining variations in latent-space separability among different kinds of models to reveal the rank correlation between our method and large-scale traditional red-teaming trends; Intra-Model Dimension (DPO Fine-Tuning Dynamics): analyzing how changes induced by DPO-based fine-tuning are reflected in our JSS proxy metric, thereby testing the sensitivity and faithfulness of JSS to alignment-related weight updates.
+Figure 5: Comparison of ranking consistency between JSS-proxy and traditional red-teaming metrics across varying numbers of models ($n$) and evaluation metrics ($k$), measured by Kendall’s $\tau$.
+4.1.1 Cross-Model Validation of JSS-Derived Safety Proxies
+For model-level differentiation, we collected a diverse set of models and evaluated their performance under a series of standardized red-teaming safety tests (20+ testing protocols with over 7000 test cases, more data listed in the Appendix table  LABEL:tab:jb_safety_split_A and  LABEL:tab:jb_safety_split_B).
+For each model, the results were aggregated into a single averaged traditional safety score.
+Our goal in this experiment is twofold: (1) to derive proxy safety indicators from our JSS-based evaluation framework, and (2) to assess the consistency between the proxy-based ranking and the ranking obtained from conventional red-teaming safety evaluations.
+Building upon the JSS evaluation indices defined earlier, we construct two higher-level proxy metrics that capture complementary aspects of safety separability.
+$\mathrm{JB/PB\ Ratio}=\frac{\sum_{s\in[0,1]}\ \mathrm{JSS}_{s}\big(\mathrm{J},\mathrm{B}\big)}{\sum_{s\in[0,1]}\ \mathrm{JSS}_{s}\big(\mathrm{P},\mathrm{B}\big)}$ (8)
+A higher value of $\mathrm{JB/PB\ Ratio}$ indicates that Jailbreak trajectories are more strongly separated from the Baseline than PlainQuery trajectories, reflecting the model’s latent ability to discriminate adversarial jailbreak behavior beyond explicit harmful intent.
+$\mathrm{JR\ Min/Max}=\frac{\min_{s\in[0,1]}\ \mathrm{JSS}_{s}\big(\mathrm{J},\mathrm{R}\big)}{\max_{s\in[0,1]}\ \mathrm{JSS}_{s}\big(\mathrm{J},\mathrm{R}\big)}$ (9)
+A higher value of $\mathrm{JR\ Min/Max}$ indicates that Jailbreak–Ideal Refusal separability remains stable across the entire trajectory, whereas a low value reflects late-stage collapse of refusal and successful jailbreak intrusion.
+We quantify the agreement between our proxy-based safety ranking $r_{\text{jss-proxy}}$ and the ranking obtained from conventional red-teaming evaluations $r_{\text{Hybrid}}$ using Kendall’s $\tau$ correlation coefficient: $\tau(r_{\text{jss-proxy}},r_{\text{Hybrid}})\in[-1,1].$
+Figure  4 show that cross a sufficiently large model set, the proxy-based ranking exhibits strong concordance with the red-teaming ground truth.
+This demonstrates that our JSS-derived proxy indicators can accurately reproduce traditional safety orderings within representative alignment regimes, while requiring far lower evaluation cost.
+As the model and metric pool expands in Figure  5, $\tau$ remains consistently high, with $p$-values far below conventional significance thresholds(<0.05<0.05).
+This indicates that even under diverse models and red-teaming metrics, the consistency between the proxy-based and red-teaming rankings remains higher than red-teaming metrics themselves and statistically significant.
+Figure 6: Evolution of JSS and behavioral safety metrics under Forward vs.
+Inverse DPO.
+The dynamic evolution of both surface-level safety behavior metrics (Unsafe Rate) and latent proxy metrics (JSS).
+The three columns respectively correspond to separability relations across different dialogue families.
+This confirms that JSS reliably tracks the direction of safety strengthening or weakening, providing a stable latent–surface correspondence.
+4.1.2 Consistency under Model Weight Variations
+In essence, we demonstrate that the geometric safety scale captured by JSS not only reflects static differences among models but also dynamically responds to behavioral transitions induced by weight adjustments.
+We introduce two opposite preference optimization directions—Forward DPO and Inverse DPO—to respectively enhance and attenuate safety alignment by fine-tuning the same base model with reversed preferences.
+Both settings use the standard DPO objective: $\mathcal{L}_{\mathrm{DPO}}(\theta)=-\,\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\Big[\log\sigma\!\big(\beta\,\Delta_{\theta}\big)\Big]$ $\text{where }\;\Delta_{\theta}=\log p_{\theta}(y^{+}\!\mid\!x)-\log p_{\theta}(y^{-}\!\mid\!x)$ In Forward DPO, safety-aligned responses are preferred over harmful ones ($y^{+}$ = safe, $y^{-}$ = harmful), whereas Inverse DPO reverses this preference to favor harmful responses ($y^{+}$ = harmful, $y^{-}$ = safe).
+To avoid conclusions driven by a single fine-tuning run or a single training step, we analyze the training-time co-evolution between a latent proxy (JSS) and surface safety metrics (UR and RR) under both Forward and Inverse DPO.
+Figure 6 shows that JSS changes track the whole training dynamics of safety behavior: when safety improves under Forward DPO, JSS increases together with decreased unsafe surface metrics; when safety is weakened under Inverse DPO, JSS decreases as surface safety degrades (Unsafe Rate increases).
+Importantly, the turning points of the dashed JSS curves appear earlier than the corresponding shifts in output behavior-based metrics (detailed information in the Appendix Table 4).
+Figure 7 further supports this temporal relation at the step level: JSS exhibits strong step-wise coupling with both UR and RR across layers (Unsafe Rate and Refusal Rate definition listed in Appendix A).
+Together, these results indicate that latent separability (JSS) changes earlier than refusal-rate changes during training, making JSS a practical early-warning signal for upcoming safety degradation, while unsafe-rate changes largely reflect the same-step training state.
+Figure 7: Step-Wise Safety Metrics (UR: Unsafe Rate; $R$: Refusual Rate) and JSS Correlation Heatmap. JSS exhibits strong step-wise coupling with both UR and RR across layers
+Attack $\mathrm{tok}_{\mathrm{in}}$ (k) $\mathrm{tok}_{\mathrm{out}}$ (k) $\mathrm{tok}$ (k)
+ArtPrompt 50.92(159) 70.10(216) 121.02(375)
+Cipher 515.65(–) 151.90(63) 667.55(63)
+DAN 502.69(–) 99.76(59) 602.45(59)
+DevMode 519.71(–) 59.74(48) 579.45(48)
+MultiLing –(–) 167.87(17) 167.87(17)
+AutoDAN 1354.89(129) 2224.48(185) 3579.37(314)
+GPTFuzzer 836.48(82) 700.46(76) 1536.94(158)
+PAIR 1212.90(123) 775.71(85) 1988.61(207)
+Mixed (PyRIT) 4797.30(–)
+Ours (JSS-only) (–) 1.03(–) 1.03(–)
+Table 1: Attack statistics aggregated over single vs.
+multi-round settings.
+Each entry is reported in thousands of tokens (k(avg)), where $\mathrm{tok}_{\mathrm{in}}$ denotes input tokens to the subject model, $\mathrm{tok}_{\mathrm{out}}$ denotes output tokens generated for evaluation, and $\mathrm{tok}$ denotes their sum.
+4.2 Baseline Comparison: Cost and Efficiency of Traditional Red-Teaming vs. Our JSS-Only Pipeline
+We compare the evaluation cost of our JSS-only pipeline with traditional online red-teaming under matched task scale and sample size.
+The baseline methods that contain single-turn, multi-turn and mixed strategy follow a standard red-teaming workflow with separate attacker, subject, and evaluator models, while our method relies solely on a several forward passes of the subject model on offline attack trajectories, without online generation or auxiliary evaluator inference.
+Specifically, we provide more details of the cost model, token accounting, and timing formulation in Appendix D when we conduct the experiment on mixed strategy (PyRIT Munoz et al.
+(2024)).
+The results support our method’s advantages as well in Appendix Figure 12.
+In table  1 we can see that across all red-teaming intensities and evaluation settings, the JSS-only pipeline consistently achieves orders-of-magnitude reductions in both token consumption.
+This efficiency gain stems from eliminating iterative generation and multi-model interaction, thereby removing the primary scalability bottleneck of traditional red-teaming.
+Figure 8: Rank correlation stability of JSS-based safety ordering under varying sample sizes for jailbreak and benign data. Appendix Fig 11 has more information about this.
+4.3 Model Sensitivity and Systemic Stability Experiments
+If our method is indeed capturing an intrinsic property of the model’s alignment dynamics—rather than dataset-specific artifacts—then its induced safety ordering should remain stable when we perturb the underlying data and sampling conditions.
+The goal of this section is therefore to test the systemic stability of our JSS-based safety scale.
+Concretely, we ask: (i) whether the relative JSS differences between models remain consistent when we vary which jailbreak trajectories are used under different attack datasets and how many sample size $n$ trajectories are included; and (ii) whether the same stability holds when we perturb the benign reference distribution, by changing the source corpus and the number of benign samples used to estimate the natural manifold.
+As shown in Fig 8 and Table 2, rank correlations consistently increase with sample size and rapidly converge to high values, indicating that the induced model ordering is not sensitive to the specific choice of attack or benign datasets.
+Lower correlations mainly arise under heterogeneous datasets or small sample sizes—particularly for jailbreak data—where weak or localized separability amplifies noise, but these effects vanish as rankings converge at moderate to large $n$.
+Source Metric Matrix Dim Mean Std Q25 Q75
+benign kendall n 0.78 0.11 0.65 0.87
+dataset 0.67 0.14 0.56 0.78
+model 0.73 0.36 0.33 1.00
+spearman n 0.88 0.08 0.79 0.94
+dataset 0.79 0.11 0.72 0.87
+model 0.76 0.34 0.40 1.00
+jailbreak kendall n 0.87 0.14 0.78 0.96
+dataset 0.45 0.24 0.33 0.60
+model 0.54 0.12 0.47 0.61
+spearman n 0.94 0.08 0.92 0.99
+dataset 0.57 0.28 0.42 0.79
+model 0.68 0.12 0.58 0.78
+Table 2: Summary of rank correlation statistics for JSS-based safety ordering.
+The Matrix Dim column indicates the axis along which rankings are compared: n denotes varying sample sizes, dataset denotes different jailbreak or benign data sources, and model denotes comparisons across model variants.
+Statistics are reported as mean, standard deviation, and interquartile range (Q25–Q75).
+A detailed pair-wise heatmap is at Appendix Fig 13.
+5 Conclusion
+This paper presents N-GLARE, a non-generative, representation-based framework for LLM safety evaluation that captures safety state through latent trajectory geometry rather than outputs.
+By modeling a benign manifold and using JSS to measure deviations of jailbreak and refusal trajectories, N-GLARE produces model-level safety rankings consistent with large-scale red-teaming across over 40 models, while reducing cost by more than two orders of magnitude.
+It further reveals fine-grained safety dynamics, demonstrating latent trajectory analysis as a scalable tool for continuous safety diagnostics.
+6 Limitations
+While N-GLARE demonstrates strong empirical effectiveness, several limitations remain.
+First, although our evaluation covers over 40 models with diverse sizes and alignment settings, it does not exhaustively represent all architectures, training paradigms, or deployment variants.
+Second, N-GLARE requires access to internal hidden representations, which limits its direct applicability to fully black-box models that expose only text outputs.
+In addition, the geometric abstraction underlying N-GLARE—modeling benign behavior as a low-dimensional manifold and inferring safety from trajectory separability—may oversimplify more complex latent structures in large models.
+Finally, although the framework avoids output generation, it still depends on the design of probing conditions, whose coverage and quality can influence the resulting safety indicators.
+The limitations above also point to several promising directions for future research.
+Model coverage can be naturally expanded as new architectures and checkpoints become available, enabling broader validation of representation-based safety indicators.
+The reliance on hidden states motivates future work on approximate or surrogate signals that could extend N-GLARE–style analysis to partially observable or API-only settings.
+Beyond harmfulness and refusal, the trajectory-based framework can be generalized to capture other alignment dimensions, such as bias, factual reliability, or long-horizon agentic safety, by defining task-specific probing conditions and reference manifolds.
+More expressive geometric models, including non-linear or multi-manifold representations, may further improve fidelity to complex latent structures.
+Finally, systematic and automated probing design offers a new research direction, positioning latent trajectory analysis as a foundation for scalable, continuous alignment monitoring and comparative safety evaluation.
+Appendix A Unsafe Rate and Refusual Rate Definition
+Unsafe Rate. This metric is computed using the safety classifications provided by Qwen3Guard. A model output is counted as unsafe whenever the guard model labels it as “Unsafe.” Formally,
+$\text{Unsafe Rate}=\frac{N_{\text{Unsafe}}}{N_{\text{Total}}},$
+where $N_{\text{Unsafe}}$ denotes the number of responses flagged as Unsafe by the guard model, and $N_{\text{Total}}$ denotes the total number of model responses evaluated.
+Refusal Rate.
+This metric is based on refusal-style keyword matching.
+A model response is counted as a refusal if it contains typical refusal patterns such as “I’m sorry”, “I cannot comply”, “I refuse to answer”, or similar formulations.
+Formally,
+$\text{Refusal Rate}=\frac{N_{\text{Refusal}}}{N_{\text{Total}}},$
+where $N_{\text{Refusal}}$ denotes the number of responses that match the refusal keyword patterns.
+Complementarity.
+These two metrics capture opposite behavioral tendencies and are therefore complementary.
+Lower Unsafe Rate indicates stronger safety, while lower Refusal Rate indicates better helpfulness and less over-defensiveness.
+A high Unsafe Rate signals safety failures, whereas a high Refusal Rate reflects an overly conservative model that rejects too many user requests.
+Appendix B Empirical Analysis of Latent Safety Trajectories
+This appendix provides structured empirical evidence and metric definitions supporting the methodological insights presented in Section 3.5.
+All observations are validated through quantitative analysis on latent representation trajectories.
+B.1 Trajectory Evolution Across Dialogue Families
+Observation.
+Model safety manifests as a dynamic process in latent representation space rather than a static output attribute.
+Protocol.
+We extract APT trajectories for four dialogue families: Baseline, PlainQuery, Jailbreak, and Ideal Refusal. Each trajectory is traced across layers and decoding steps under identical model settings.
+Empirical Evidence.
+Figure 3 shows that:
+  • Baseline trajectories remain flat and low-variance, forming a stable reference manifold.
+  • PlainQuery trajectories diverge sharply from Baseline across all layers, indicating rapid risk detection and refusal activation.
+  • Jailbreak trajectories gradually deviate from Baseline and exhibit downward drift in higher layers, reflecting progressive erosion of refusal mechanisms.
+  • Ideal Refusal initially overlaps with Jailbreak but later bifurcates, indicating a latent distinction between compliant and defensive generation paths.
+Takeaway.
+Safety-relevant information is encoded in the relative geometry of trajectories rather than in isolated representations.
+B.2 Jailbreak–Baseline Separability and Detection Capability
+Observation.
+Separability between Jailbreak and Baseline trajectories reflects the model’s latent capability to detect jailbreak process shifting.
+Metric and Protocol.
+To test whether Jailbreak trajectories encode stage-dependent information, we partition the jailbreak process into early, mid, and late stages.
+A lightweight neural classifier is trained to distinguish representations from different stages.
+If stage-wise structure is present, classification accuracy should exceed chance level.
+Empirical Evidence.
+For aligned LLaMA-family models, stage classification accuracy reaches:
+Table 3: Stage-wise classification accuracy for Jailbreak trajectory representations in aligned LLaMA-family models.
+Stage Pair Classification Accuracy
+Early vs. Mid 0.86 – 0.97
+Early vs. Late 0.91 – 0.96
+Mid vs. Late 0.85 – 0.86
+Higher-than-chance accuracy across all stage pairs indicates a reliable and alignment-dependent phase structure in latent jailbreak trajectories.
+After ablating safety alignment (manually jailbroken variants), accuracy collapses to approximately $0.5$ for all pairs.
+Takeaway.
+Jailbreak trajectories exhibit a reliable, alignment-dependent phase structure, supporting the use of Jailbreak–Baseline separability as a proxy for latent detection capability.
+B.3 Jailbreak–Ideal Refusal Separability and Refusal Tendency
+Observation.
+Separability between Jailbreak and Ideal Refusal trajectories reflects intrinsic refusal tendency.
+Metric Definition (ANM).
+We quantify refusal inclination using the Activation-based Negation Measure (ANM):
+$\mathrm{ANM}(x)=\sum_{v\in\mathcal{V}}W_{\mathrm{ref}}(v)\,p(v\mid x),$ (10)
+where $p(v\mid x)$ is the next-token distribution and $W_{\mathrm{ref}}$ is a fixed refusal-prototype distribution constructed from normalized embeddings of refusal seed phrases.
+Empirical Evidence.
+Across layers, ANM curves exhibit a consistent mid-stage rise followed by late-stage collapse.
+This pattern aligns with the Ideal Refusal–Jailbreak JSS waveform.
+Layer-wise correlations between ANM and Ideal Refusal–Jailbreak JSS are:
+$r=0.70,\ 0.56,\ 0.54\quad(\text{RMSE}=0.13\text{--}0.20).$
+Figure  9 also shows that mid-stage elevation corresponds to increased refusal probability, while late-stage decline coincides with jailbreak success.
+Takeaway.
+Latent trajectory separability is tightly coupled with explicit refusal behavior, validating its use as an interpretable proxy for refusal stability.
+Figure 9: Coupled dynamics of JSS and ANM along the jailbreak process.
+The color indicates the normalized refusal-intensity score (ANM); The inset line plot in the upper-right overlays slice-wise JSS and ANM curves over the standardized progress axis, revealing a shared mid-phase rise and late-phase collapse, which highlights the strong correlation between geometric separability (JSS) and probability-level refusal tendency (ANM).
+Appendix C Construction of the Benign Manifold and Geometric Turning Angle
+This appendix provides the full mathematical formulation underlying the benign manifold, the associated normal direction, and the geometric turning angle introduced in Section 3.3.
+These derivations are omitted from the main text for clarity and readability.
+C.1 Benign Manifold Construction
+To characterize normal generation behavior, we construct a benign manifold using latent representations collected under benign inputs.
 For each group $G$, we aggregate all benign activations across prompts and progress values:
 $\mathcal{H}_{B}^{(G)}=\{h^{(B,G)}\}.$ (11)
-We estimate the empirical mean and covariance
-$\begin{split}\mu^{(G)}&=\mathbb{E}[h^{(B,G)}],\qquad\\ \Sigma^{(G)}&=\mathbb{E}\bigl[(h^{(B,G)}-\mu^{(G)})(h^{(B,G)}-\mu^{(G)})^{\top}\bigr],\end{split}$ (8)
-and perform a rank-$r$ eigendecomposition
-$\Sigma^{(G)}\approx U_{r}^{(G)}\Lambda_{r}^{(G)}{U_{r}^{(G)}}^{\top},\quad r\ll d,$ (9)
-where $U_{r}^{(G)}\in\mathbb{R}^{d\times r}$ contains the top-$r$ eigenvectors and $\Lambda_{r}^{(G)}\in\mathbb{R}^{r\times r}$ is the diagonal matrix of corresponding eigenvalues.
-This defines a benign subspace
-$\mathcal{S}^{(G)}=\mathrm{span}\bigl(U_{r}^{(G)}\bigr),$ (10)
-and a whitening transform
-$W^{(G)}={\Lambda_{r}^{(G)}}^{-1/2}{U_{r}^{(G)}}^{\top}.$ (11)
-For any representation $h$, its whitened coordinate is
-$z=W^{(G)}(h-\mu^{(G)}),$ (12)
-which removes scale and correlation effects in the benign distribution and provides a canonical coordinate system for measuring deviations across models.
-By the Eckart–Young–Mirsky theorem [EckartYoung1936, GolubVanLoan2013] and the tangent-space linearization principle for smooth manifolds [DoCarmo1976], the top-$r$ PCA subspace offers the statistically optimal linear surrogate of the benign manifold.
-Given the benign manifold, we define an instantaneous geometric measure that characterizes how a trajectory moves relative to this manifold.
-For any representation $h$, we consider its reconstruction from the benign subspace:
-$\hat{h}^{(G)}=\mu^{(G)}+U_{r}^{(G)}{U_{r}^{(G)}}^{\top}(h-\mu^{(G)}),$ (13)
-and define the residual
-$r^{(G)}(h)=h-\hat{h}^{(G)}=h-\mu^{(G)}-U_{r}^{(G)}{U_{r}^{(G)}}^{\top}(h-\mu^{(G)}).$ (14)
-The squared norm
-$E_{b}^{(G)}(h)=\|r^{(G)}(h)\|_{2}^{2}$ (15)
-acts as a deviation energy from the benign manifold. Its gradient
-$n^{(G)}(h)=\nabla E_{b}^{(G)}(h)=2r^{(G)}(h)$ (16)
-defines an outward normal direction pointing away from the benign subspace at $h$.
-For each trajectory and its representation $h_{t}\in\Gamma_{M}^{(A,G)}$, the tangent direction at node $t$ is given by
-$\tau_{t}^{(A,G)}=\frac{h_{t+1}^{(A,G)}-h_{t}^{(A,G)}}{\|h_{t+1}^{(A,G)}-h_{t}^{(A,G)}\|_{2}},$ (17)
-for $1\leq t<T_{A}$. We then define the turning angle $\theta_{t}^{(A,G)}\in[0,\pi]$ between the trajectory tangent and the outward normal:
+We estimate the empirical mean and covariance:
+$\begin{split}\mu^{(G)}&=\mathbb{E}[h^{(B,G)}],\\ \Sigma^{(G)}&=\mathbb{E}\bigl[(h^{(B,G)}-\mu^{(G)})(h^{(B,G)}-\mu^{(G)})^{\top}\bigr].\end{split}$ (12)
+A rank-$r$ eigendecomposition is then performed:
+$\Sigma^{(G)}\approx U_{r}^{(G)}\Lambda_{r}^{(G)}{U_{r}^{(G)}}^{\top},\qquad r\ll d,$ (13)
+where $U_{r}^{(G)}$ contains the top-$r$ eigenvectors and $\Lambda_{r}^{(G)}$ the corresponding eigenvalues. This defines a linear surrogate of the benign manifold:
+$\mathcal{S}^{(G)}=\mathrm{span}\bigl(U_{r}^{(G)}\bigr).$ (14)
+By the Eckart–Young–Mirsky theorem Eckart and Young (1936); Golub and Van Loan (2013), this subspace provides the optimal rank-$r$ approximation of the benign distribution in the least-squares sense.
+Under standard smoothness assumptions, it also serves as a local tangent-space approximation of the underlying benign manifold do Carmo (1976).
+C.2 Residual Energy and Outward Normal
+Given a representation $h$, its reconstruction from the benign subspace is:
+$\hat{h}^{(G)}=\mu^{(G)}+U_{r}^{(G)}{U_{r}^{(G)}}^{\top}(h-\mu^{(G)}).$ (15)
+The residual vector
+$r^{(G)}(h)=h-\hat{h}^{(G)}$ (16)
+captures the component orthogonal to the benign manifold. We define the residual energy:
+$E_{b}^{(G)}(h)=\|r^{(G)}(h)\|_{2}^{2}.$ (17)
+The gradient of this energy with respect to $h$ is:
+$n^{(G)}(h)=\nabla E_{b}^{(G)}(h)=2\,r^{(G)}(h),$ (18)
+which defines an outward normal direction pointing away from the benign manifold. Up to a constant factor, the residual vector itself therefore serves as the normal direction used in practice.
+C.3 Trajectory Tangent Direction
+For a representation trajectory $\Gamma_{M}^{(A,G)}=\{h_{t}^{(A,G)}\}_{t=1}^{T_{A}}$, the local tangent direction at step $t$ is defined as:
+$\tau_{t}^{(A,G)}=\frac{h_{t+1}^{(A,G)}-h_{t}^{(A,G)}}{\|h_{t+1}^{(A,G)}-h_{t}^{(A,G)}\|_{2}},\qquad 1\leq t<T_{A}.$ (19)
+This tangent captures the instantaneous direction of representation evolution along the inference process.
+C.4 Derivation of the Geometric Turning Angle
+At each step $t$, we measure how the trajectory direction aligns with the outward normal of the benign manifold. The geometric turning angle is defined as:
 $\theta_{t}^{(A,G)}=\arccos\left(\frac{\tau_{t}^{(A,G)\top}n^{(G)}\bigl(h_{t}^{(A,G)}\bigr)}{\|\tau_{t}^{(A,G)}\|_{2}\,\|n^{(G)}\bigl(h_{t}^{(A,G)}\bigr)\|_{2}}\right).$ (20)
-In practice, since $n^{(G)}(h)$ is collinear with $r^{(G)}(h)$, we compute the numerator as $\tau_{t}^{(A,G)\top}r^{(G)}\bigl(h_{t}^{(A,G)}\bigr)$ and normalize accordingly.
-Intuitively, $\theta_{t}^{(A,G)}$ measures how strongly the trajectory at $t$ turns “away” from the benign manifold.
-Smaller angles (particularly $\theta_{t}^{(A,G)}<\pi/2$) indicate movement into regions further from benign behavior.
-Algorithm 1 JSS Core Computation for Model $M$
+Since $n^{(G)}(h)$ is collinear with $r^{(G)}(h)$, the angle can equivalently be computed using the residual direction.
+By construction, $\theta_{t}^{(A,G)}\in[0,\pi]$.
+Smaller angles indicate that the trajectory is moving more directly away from the benign manifold, corresponding to stronger deviation from normal generation behavior.
+This turning angle serves as the local geometric primitive for trajectory-level aggregation in the main text.
+Appendix D Cost Model and Efficiency Analysis
+This appendix details the cost modeling assumptions, token accounting rules, and runtime decomposition used in the efficiency comparison between the traditional online red-teaming pipeline and our JSS-only evaluation pipeline.
+All measurements are conducted under matched task scale and sample size to ensure a fair comparison.
+D.1 Evaluation Pipelines and Model Roles
+We compare two evaluation pipelines:
+JSS-only pipeline (Ours).
+Our method operates on offline multi-turn attack trajectories.
+For each example, we perform a single forward pass on the subject model $M_{s}$ to extract hidden representations.
+No online generation, attack synthesis, or evaluator inference is required.
+Traditional red-teaming pipeline (Baseline).
+The baseline follows a standard online red-teaming workflow implemented using PyRIT with Ollama backends, involving three model roles:
+  • $M_{r}$: red-team attack generator,
+  • $M_{s}$: subject model under evaluation,
+  • $M_{e}$: evaluator or safety discriminator.
+Each attack session proceeds in multiple turns, with all three models invoked at every turn.
+D.2 Token Cost Formulation
+Let $N$ denote the number of attack targets, $\bar{T}_{r}$ the average number of attack turns per session, and $\bar{L}_{r}$, $\bar{L}_{s}$, $\bar{L}_{e}$ the average token lengths generated by $M_{r}$, $M_{s}$, and $M_{e}$, respectively.
+Baseline.
+The total generated-token cost of traditional red-teaming is:
+$C_{\text{token}}^{\text{Base}}(N)=N\cdot\bar{T}_{r}\cdot\bigl(\bar{L}_{r}+\bar{L}_{s}+\bar{L}_{e}\bigr).$ (21)
+JSS-only pipeline.
+Our method requires only a forward pass over an offline multi-turn dataset. For bookkeeping consistency, we count one subject-model forward pass per turn as one token-equivalent unit:
+$C_{\text{token}}^{\text{Ours}}(N)=N\cdot\bar{T}_{\text{offline}},$ (22)
+where $\bar{T}_{\text{offline}}$ denotes the average number of turns in the offline attack trajectories.
+D.3 Runtime Cost Decomposition
+Baseline.
+The total runtime cost of the traditional pipeline decomposes as:
+$C_{\text{time}}^{\text{Base}}=\mathrm{Time}_{M_{r}}+\mathrm{Time}_{M_{s}}+\mathrm{Time}_{M_{e}},$ (23)
+where each term includes both model inference time and associated system overheads incurred at every attack turn.
+JSS-only pipeline.
+The runtime cost of our method is:
+$C_{\text{time}}^{\text{Ours}}=\mathrm{Time}_{M_{s}^{\text{forward}}}+\mathrm{Time}_{\text{offline\_JSS}}+\mathrm{Time}_{\text{benign\_con}}.$ (24)
+Here, $\mathrm{Time}_{M_{s}^{\text{forward}}}$ denotes the time required to perform forward passes on the subject model over the offline dataset.
+The costs of benign distribution construction and JSS computation are typically negligible compared to model inference and are therefore omitted in efficiency comparisons.
+D.4 Measurement Protocol
+We evaluate both pipelines at multiple red-teaming intensities $N\in\{n_{1},n_{2},\dots\}$.
+For the baseline, online attacks are generated using PyRIT, and all generated tokens from $M_{r}$, $M_{s}$, and $M_{e}$ are counted. Wall-clock time is measured end-to-end for each evaluation run.
+For the JSS-only pipeline, the same set of attack instructions is used in the form of offline multi-turn datasets derived from RefEx and PlainR.
+We record token-equivalent counts and elapsed wall-clock time for the forward pass and offline analysis.
+D.5 Evaluation Metrics
+We compare the two pipelines using the following efficiency metrics:
+  • Total generated-token count: overall system token usage.
+  • Subject-model token share: fraction of tokens attributed to the evaluated model $M_{s}$.
+  • Per-token time cost: wall-clock time normalized by total generated tokens, serving as a throughput indicator.
+Lower per-token time cost indicates higher evaluation efficiency.
+D.6 Summary of Efficiency Results
+Across all red-teaming intensities and evaluation settings, the JSS-only pipeline consistently reduces both token consumption and runtime by orders of magnitude compared to the traditional online red-teaming pipeline.
+This reduction arises from eliminating iterative generation and multi-model interaction, thereby removing the primary scalability bottleneck of conventional red-teaming approaches.
+Algorithm 1 JSS Core Computation for Model $M$, For clarity, the pseudocode illustrates a single run on the full dataset.
+In our experiments, we wrap this procedure in a bootstrap loop to obtain empirical confidence intervals and perform non-parametric statistical tests across models.
 0: Model $M$; datasets $\mathcal{D}_{A}$ for $A\in\{B,J,R,P\}$; hyperparameters $(r,G,I,s_{0},\{w_{i}\},\{w_{i}^{\prime}\})$.
@@
 19:    $\mathrm{JSS}^{(G)}(A,B^{\prime})\mathrel{+}=\mathrm{JS}_{i}^{(G)}(A,B^{\prime})$
-20:    if $s_{i}\leq s_{0}$ then
-21:     $\mathrm{JSS}^{(G)}_{\text{early}}(A,B^{\prime})\mathrel{+}=\mathrm{JS}_{i}^{(G)}(A,B^{\prime})$
-22:    end if
+20:   end for
+21:  end for
+22:  $\mathrm{JSS}(A,B^{\prime})\leftarrow\frac{1}{|G|}\sum_{G}\mathrm{JSS}^{(G)}(A,B^{\prime})$
 23:end for
-24:  end for
-25:  $\mathrm{JSS}(A,B^{\prime})\leftarrow\frac{1}{|G|}\sum_{G}\mathrm{JSS}^{(G)}(A,B^{\prime})$
-26:  $\mathrm{JSS}_{\text{early}}(A,B^{\prime})\leftarrow\frac{1}{|G|}\sum_{G}\mathrm{JSS}^{(G)}_{\text{early}}(A,B^{\prime})$
-27:end for
-28:return $\mathrm{JSS}(J,B)$, $\mathrm{JSS}_{\text{early}}(J,B)$, $\mathrm{JSS}(J,R)$, $\mathrm{JSS}_{\text{early}}(J,R)$, $\mathrm{JSS}(P,B)$
-4.3 Probing Conditions, Slicing, and JSS Metrics
-We apply the above constructions to three probing conditions that are designed to be isomorphic along the progress axis $s$:
-  • Benign ($B$): natural or conventional task distributions, used to fit the benign manifold and serve as a reference.
-  • Jailbreak ($J$): adversarial sequences designed to elicit harmful behavior.
-  • Ideal refusal ($R$): sequences identical to $J$ in user content, but with the assistant response replaced by a fixed, policy-compliant refusal template.
-  • Plain Query ($P$): sequences that repeatedly present the raw jailbreak true intention to the model, capturing the model’s unconstrained internal response trajectory under direct exposure to harmful prompts.
-For each group $G$ and condition $A$, we thus obtain a set of angle–progress pairs (i.e. Angular-Probabilistic Trajectories, APT)
-$\Theta^{(A,G)}=\{(\theta_{t}^{(A,G)},s_{t})\}.$ (19)
-To study how separability evolves along the interaction, we discretize the standardized axis $s\in[0,1]$ into $I$ slices (bins) $\{s_{i}\}_{i=1}^{I}$.
-In slice $s_{i}$, we collect all angles whose progress falls into that interval:
-$\Theta_{A}^{(G)}(s_{i})=\{\theta_{t}^{(A,G)}:s_{t}\in s_{i}\}.$ (20)
-We also define an early-stage threshold $s_{0}\in(0,1)$ (default $s_{0}=0.4$) and use it to construct “early-ss” metrics that focus on whether the model exhibits safe behavior in the initial portion of the trajectory.
-Within each slice $s_{i}$ and group $G$, we estimate the empirical distributions of the angle samples for two conditions $A$ and $B$ (e.g., $J$ vs.
-$B$, or $J$ vs.
-$R$), denoted by $\mathrm{Dist}(\Theta_{A}^{(G)}(s_{i}))$ and $\mathrm{Dist}(\Theta_{B}^{(G)}(s_{i}))$, meaninig the distribution of them.
-We then compute the Jensen–Shannon divergence:
-$\begin{split}\mathrm{JS}_{i}^{(G)}(A,B)=\mathrm{JS}\bigl(\mathrm{Dist}(\Theta_{A}^{(G)}(s_{i}))\,\big\|\,\mathrm{Dist}(\Theta_{B}^{(G)}(s_{i}))\bigr)\end{split}$ (21)
-To obtain a model-level separability measure, we aggregate slice-level divergences. For each group $G$,
-$\mathrm{JSS}^{(G)}(A,B)=\sum_{i=1}^{I}\,\mathrm{JS}_{i}^{(G)}(A,B),$ (22)
-and average across layer groups to obtain the all-s separability:
-$\mathrm{JSS}(A,B)=\frac{1}{|G|}\sum_{G}\mathrm{JSS}^{(G)}(A,B).$ (23)
-Restricting to slices with $i\leq s_{0}$ yields the early-s separability:
-$\mathrm{JSS}_{\text{early}}(A,B)=\frac{1}{|G|}\sum_{G}\sum_{i\leq s_{0}}\mathrm{JS}_{i}^{(G)}(A,B).$ (24)
-Algorithm 1 summarizes the core computation of JSS metrics for a single model $M$.
-For clarity, the pseudocode illustrates a single run on the full dataset.
-In our experiments, we wrap this procedure in a bootstrap loop to obtain empirical confidence intervals and perform non-parametric statistical tests across models.
-(a) Lower Group
-(b) Middle Group
-(c) Upper Group
-(d) Lower Group
-(e) Middle Group
-(f) Upper Group
-Figure 2: APT trajectories collected from different layer groups.
-Subfigures (a)(b)(c) correspond to safety-aligned models, while (d)(e)(f) are from jailbreak-tuned models.
-Red curves represent Plain Query trajectories, yellow denote Benign trajectories, blue indicate Jailbreak trajectories, and green correspond to Ideal Refusal trajectories.
-Safety-aligned LLMs exhibit well-separated and low-variance APT trajectories, whereas jailbreak-tuned LLMs show weakened separability and substantially higher variance across all query types, indicating degraded safety boundaries.
-5 Experiments and Analysis
-This section begins by presenting the intuitive safety characteristics derived from latent representation trajectories, illustrating how model representations evolve under different dialogue families.
-Subsequently, from a dynamical perspective, we systematically examine the correspondence between these latent features and the model’s surface-level safety behaviors, thereby validating the safety representation hypothesis proposed earlier.
-On this basis, we employ quantified safety feature metrics to evaluate various aligned models and compare their rankings with those obtained from systematic traditional red-teaming evaluations, demonstrating a high degree of consistency in safety scales as well as a significant advantage in evaluation cost.
-Finally, we investigate the statistical consistency of the ranking results across different datasets and sample sizes, analyzing the stability of this evaluation framework under changes in data distribution and sampling conditions.
-5.1 Trajectory Evolution Dynamics
-This subsection empirically establishes one core observation of our evaluation framework:
-Observation 1.
-A model’s safety state is not only a behavior-based static attribute, but rather a dynamic process that can be continuously tracked within the latent representation space.
-Specifically, in Figure 2, we compare the APT evolutionary trajectories of the PlainQuery, Baseline, Jailbreak, and Ideal Refusal dialogue families.
-Baseline maintains a stable, flat trajectory, representing the model’s consistent generation direction during natural dialogue.
-PlainQuery rises sharply across all layers, rapidly diverging from the baseline, indicating that the model promptly shifts its generation direction upon detecting harmful intent, forming a distinct refusal response.
-Jailbreak trajectories gradually diverge upward from the baseline, exhibiting a notable downward drift in the higher layers, suggesting that adversarial semantics accumulate in upper-layer representations, progressively eroding the refusal mechanism.
-Although the Ideal-Refusal trajectory initially follows the Jailbreak path, a clear bifurcation emerges, indicating that the model encodes a latent distinction between accepting and refusing harmful instructions.
-Such separation suggests that the model possesses an intrinsic cognitive bias that differentiates unsafe compliance from defensive rejection within its internal representation space.
+24:return $\mathrm{JSS}(J,B)$, $\mathrm{JSS}(J,R)$, $\mathrm{JSS}(P,B)$
+Table 4: Temporal lag between JSS changes and Forward DPO surface-level metrics across dialogue families (UR: Unsafe Rate; $R$: Refusal Rate).
+Lag Prev.
+denotes the percentage of trajectories in which JSS leads the corresponding surface metric, and Median Shift reports the median temporal offset (in training steps), with negative values indicating that changes in JSS precede changes in the surface metric.
+Metric Family Lag Prev. Median Shift
+JSS vs. RR RefEx–JB 74 4.4-4.4
+JSS vs. RR JB–Baseline 64 1.4-1.4
+JSS vs. UR JB–Baseline 53 0.6-0.6
 (a) Qwen Family
@@
 Figure 10: Multi-perspective cross-model safety ranking consistency between N-GLARE and traditional red-teaming.
-For each model family ((a–f; see Appendix Table II for model abbreviations on scatters)), the left scatter plot compares the relative ranks induced by our JSS-based mixed proxy (X-axis) and by Hybrid red-teaming scores (Y-axis); points lying near the diagonal blue band indicate close agreement between the two rankings.
+For each model family ((a–f; see Appendix Table 5 for model abbreviations on scatters)), the left scatter plot compares the relative ranks induced by our JSS-based mixed proxy (X-axis) and by Hybrid red-teaming scores (Y-axis); points lying near the diagonal blue band indicate close agreement between the two rankings.
 The right panel in each family shows the corresponding raw JSS-proxy (blue) and Hybrid (red) scores over all subset models, with shaded regions denoting bootstrap variability, again yielding almost identical within-family orderings.
-Note that ”Hybrid” is the overall average across score all tested red-teaming tasks, with detailed task-wise scores provided in Appendix Table LABEL:tab:jb_safety_multi.
-5.2 Synchronous Evolution of Jailbreak Progression, Refusal Tendency, and JSS Dynamics
-After obtaining the intuitive empirical phenomena, we focus on verifying two key observations:
-Observation 2.
-Divergence between Jailbreak and Baseline trajectories as an indicator of the model’s discriminative capability to detect jailbreak attacks process shifting.
-A separation between the jailbreak attack trajectory and the baseline behavior trajectory in the representation space signifies a clear process shift in the model’s internal dynamics from normal generation to adversarial alignment.
-This divergence reflects the model’s discriminative capability to detect and respond to emerging jailbreak cues, serving as a direct and quantifiable indicator of its latent safety awareness.
-Observation 3.
-Distance between Jailbreak and Ideal Refusal trajectories as an indicator of the model’s intrinsic refusal tendency.
-The JS distance between Jailbreak and Ideal Refusal trajectories directly reflects the model’s ability to distinguish and resist jailbreak behaviors.
-An increase in this distance indicates that the model has a stronger latent discrimination between “accepting” and “blocking” jailbreak instructions, thus exhibiting an enhanced intrinsic refusal tendency, corresponding to a higher level of safety awareness.
-5.2.1 Phase-wise Dynamics of Jailbreak Trajectories and Latent Safety Awareness
-For observation 2, if a model lacked latent safety awareness, JB activations would remain near the Baseline manifold, yielding no meaningful temporal segmentation.
-To test whether phase structure exists and whether it is safety-dependent, we use a lightweight neural classifier as an operational probe: if representations at different jailbreak stages encode distinct information, the classifier should reliably separate them.
-Aligned Llama-family models indeed show strong separability: 0.86–0.97 (start vs.
-mid), 0.91–0.96 (start vs.
-end), 0.85–0.86 (mid vs.
-end).
-However, after ablating safety alignment (manually jailbroken variants), stage classification collapses to $\sim 0.33$ for all pairs.
-This contrast provides two conclusions: (1) the phase structure is statistically reliable, because it can be consistently decoded by a simple classifier; (2) it is safety-dependent, because the structure disappears once alignment is destroyed.
-Thus, aligned models internally encode a latent variable corresponding to the stage of the jailbreak process, and the resulting phase-wise divergence of JB from Baseline offers a direct, graded indicator of latent safety awareness.
-Figure 4: Coupled dynamics of JSS and ANM along the jailbreak process.
-The color indicates the normalized refusal-intensity score (ANM); The inset line plot in the upper-right overlays slice-wise JSS and ANM curves over the standardized progress axis, revealing a shared mid-phase rise and late-phase collapse, which highlights the strong correlation between geometric separability (JSS) and probability-level refusal tendency (ANM).
-5.2.2 Distance between Jailbreak and Ideal Refusal Trajectories and the Model’s Refusal Tendency
-For another observation 3, empirical measurements confirm a strong synchrony between Ideal Refusal-Jailbreak separability and model’s refusal tendency.
-We evaluate the model’s latent refusal inclination strength by using a unified metric of how strongly the model tends to generate reject-related tokens at each stage of the jailbreak process.
-Specifically, we formalize ANM, which quantifies probability-level concentration on “refusal-cluster” tokens.
-$\mathrm{ANM}(x)=\big\langle W_{\mathrm{ref}},\,p(\cdot\mid x)\big\rangle=\sum_{v\in\mathcal{V}}W_{\mathrm{ref}}(v)\,p(v\mid x),$ (25)
-where $p(v\mid x)=\mathrm{softmax}(z_{v}(x)/\tau_{\text{pred}})$ is the model’s next-token distribution, and $W_{\mathrm{ref}}$ is a fixed refusal–prototype distribution constructed by:
-$W_{\mathrm{ref}}(v)=\frac{\exp(\langle\hat{e}_{v},\hat{c}\rangle/\tau_{\text{ref}})}{\sum_{u\in\mathcal{V}}\exp(\langle\hat{e}_{u},\hat{c}\rangle/\tau_{\text{ref}})},$ (26)
-with $\hat{c}$ obtained by averaging normalized embeddings of several refusal-seed phrases.
-(”sorry”, ”I can’t”, etc.) Thus, ANM quantifies how strongly the model’s predictive distribution aligns with refusal-semantics in latent space.
-Across layers, ANM curves exhibit a striking “mid-phase rise and end-phase collapse” pattern that matches the Ideal Refusal-Jailbreak JSS waveform projected into the same 0–1 domain.
-According to the results, layer-wise correlations between ANM and Ideal Refusal-Jailbreak JSS (r=0.70,0.56,0.54;RMSE=0.130.20r=0.70,0.56,0.54;\ $\text${RMSE}=0.13–0.20)[hastie2009elements] reveal a strong phase-locked coupling between representational separability and refusal intensity.
-In each layer, the mid-segment elevation ($s\approx 45–65$) corresponds to a surge of refusal semantics, while the late-stage decline ($s>70$) aligns with the erosion of this tendency as jailbreak success emerges.
-The consistent alignment between ANM and Ideal Refusal-Jailbreak separability provides a quantitative bridge between latent-space dynamics and explicit refusal behavior, establishing an interpretable and measurable indicator of the model’s latent safety awareness.
-5.2.3 Conclusion
-Overall, we demonstrate that a model’s safety is manifested not only in its observable output content, but more fundamentally in the organization and evolution of its latent representation trajectories under our method.
-Therefore, analyzing these trajectory’s separability can equally reveal the core safety mechanisms that traditional output-based evaluations seek to capture.
-5.3 Model Differentiation and Safety Characterization
-This section presents the core capability of our evaluation framework: its ability to precisely and transferably characterize safety differences across models.
-In other words, the JSS-based system not only captures the internal representational dynamics of a model but also enables multi-scale and fine-grained quantification of safety variation and underlying regularities across different alignment levels.
-Within this overarching framework, our analyses proceed along two complementary axes:
-  • Inter-Model Dimension (Model-Level Differentiation): examining subtle variations in latent-space separability among instruction-aligned models to reveal the continuous distribution of intermediate safety regimes;
-  • Intra-Model Dimension (Weight-Level Evolution): contrasting forward DPO and inverse DPO settings to observe how safety alignment evolves directionally with changes in alignment weight.
-Together, these two perspectives form a progressive precision hierarchy: from inter-model to intra-model analysis—demonstrating that the JSS framework consistently reflects alignment strength across granularities, achieving a unified, continuous, and model-invariant scale for safety evaluation.
-Figure 5: Layered subset rank consistency with statistically significant (bottom right) agreement across all $\tau$ thresholds.
-5.3.1 Cross-Model Validation of JSS-Derived Safety Proxies
-For model-level differentiation, we collected a diverse set of models and evaluated their performance under a series of standardized red-teaming safety tests (21 testing protocols with over 7000 test cases).
-For each model, the results were aggregated into a single averaged safety score (Hybrid Score).
-Our goal in this experiment is twofold: (1) to derive proxy safety indicators from our JSS-based evaluation framework, and (2) to assess the consistency between the proxy-based ranking and the ranking obtained from conventional red-teaming safety evaluations.
-Building upon the JSS evaluation indices defined earlier, we construct two higher-level proxy metrics that capture complementary aspects of safety separability.
-$\mathrm{JB/PB\ Ratio}=\frac{\sum_{s\in[0,1]}\ \mathrm{JSS}_{s}\big(\mathrm{J},\mathrm{B}\big)}{\sum_{s\in[0,1]}\ \mathrm{JSS}_{s}\big(\mathrm{P},\mathrm{B}\big)}$ (27)
-A higher value indicates that the overall separation strength between Jailbreak and Baseline trajectories is significant based on the metric that between PlainQuery and Baseline.
-This reflects stronger discriminability of jailbreak trajectories.
-$\mathrm{JR\ Min/Max}=\frac{\min_{s\in[0,1]}\ \mathrm{JSS}_{s}\big(\mathrm{J},\mathrm{R}\big)}{\max_{s\in[0,1]}\ \mathrm{JSS}_{s}\big(\mathrm{J},\mathrm{R}\big)}$ (28)
-A higher ratio indicates that the Ideal Refusal-Jialbreak separability remains stable across the entire trajectory.
-If the separation persists throughout decoding, the minimum approaches the maximum, resulting in a higher ratio.
-Conversely, if the separation collapses near the end—implying successful jailbreak intrusion and erosion of refusal—this ratio drops sharply, capturing the characteristic “mid-phase emergence–end-phase collapse” dynamics.
-To obtain a unified scalar JSS-proxy for safety alignment evaluation, we combine the two secondary metrics with equal weighting:
-$\mathrm{JSS\text{-}Proxy}=0.5*\mathrm{JB/PB\ Ratio}+0.5*\mathrm{JR\ Min/Max}.$ (29)
-We quantify the agreement between our proxy-based safety ranking $r_{\text{jss-proxy}}$ and the ranking obtained from conventional red-teaming evaluations $r_{\text{Hybrid}}$ using Kendall’s $\tau$ correlation coefficient:
-$\tau(r_{\text{jss-proxy}},r_{\text{Hybrid}})\in[-1,1].$ (30)
-Figure  5 and  3 show that cross a sufficiently large model set, the proxy-based ranking exhibits strong concordance with the red-teaming ground truth.
-This demonstrates that our JSS-derived proxy indicators can accurately reproduce traditional safety orderings within representative alignment regimes, while requiring far lower evaluation cost.
-As the model pool expands, $\tau$ shows a mild decrease but remains consistently high, with $p$-values far below conventional significance thresholds.
-This indicates that even under broader model coverage, the alignment between the proxy-based and red-teaming rankings remains statistically significant.
-Such behavior illustrates the robust hierarchical consistency of our proxy system, maintaining statistically significant agreement at scale while achieving near-perfect correspondence within focused model subsets.
-(a) Inverse DPO Baseline-Jailbreak
-(b) Inverse DPO Jailbreak-Plian refusual
-(c) Inverse DPO Jialbreak-Refuse existing
-(d) Forward DPO Baseline-Jailbreak
-(e) Forward DPO Jailbreak-Plian refusual
-(f) Forward DPO Jialbreak-Refuse existing
-Figure 6: Evolution of JSS and behavioral safety metrics under Forward vs.
-Inverse DPO.
-The dynamic evolution of both surface-level safety behavior metrics (solid lines: unsafe rate and refusal rate) and latent proxy metrics (dashed lines: JSS ) as functions of the training global step during the Forward DPO (lower row) and Inverse DPO (upper row) processes.
-The three columns respectively correspond to separability relations across different dialogue families.
-5.3.2 Consistency under Model Weight Variations
-In essence, we demonstrate that the geometric safety scale captured by JSS not only reflects static differences among models but also dynamically responds to behavioral transitions induced by weight adjustments.
-Specifically, we introduce two opposite preference directions: Forward DPO and Inverse DPO, which respectively correspond to the enhancement and attenuation of safety alignment.
-We employ the preference optimization loss to fine-tune the same base model in two opposite directions:
-$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\,\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\Big[\log\sigma\!\big(\beta\,\Delta_{\theta}\big)\Big]$ (31)
-$\text{where }\;\Delta_{\theta}=\log p_{\theta}(y^{+}\!\mid\!x)-\log p_{\theta}(y^{-}\!\mid\!x).$ Specifically, we deploy it for two purposes:
-Forward DPO. The model is fine-tuned with safety-oriented preferences, where compliant responses are favored over harmful ones ($y^{+}$ = safe, $y^{-}$ = harmful).
-Inverse DPO. We reverse the preference direction so that harmful responses are favored over compliant ones ($y^{+}$ = harmful, $y^{-}$ = safe).
-To avoid conclusions biased by a single fine-tuning intensity, we systematically generate a structured set of model variants within the LoRA parameter space, fixing the base model and dataset while varying: LoRA target modules and LoRA Rank.
-Figure  6 illustrates that from a correlation perspective, the local peaks and troughs of the JSS curves strongly align with fluctuations in the surface safety metrics, including both Unsafe Rate and Refusal Rate (Definition in Appendix Section B).
-This observation indicates that the degree of separability in latent space evolves in synchrony with the model’s explicit refusal behaviors in output space.
-A closer examination reveals a consistent time lag of several training windows between the changes in JSS and the corresponding shifts in surface metrics, with the effect being particularly evident in the early stages of Inverse DPO.
-Quantitatively, the lag is strongest under Inverse DPO (about 76% of runs exhibit lag).
-For Forward DPO, in the RefEx–JB family, 74% of trajectories show “JSS leads, refusal lags,” with a median shift of –4.4 steps, whereas in JB–Baseline the effect weakens but persists (64%, median –1.4).
-For the unsafe rate, both families show near-synchronous behavior (RefEx–JB: 45% lead / 48% lag; JB–Baseline centered around zero with median –0.6).
-Together, these results show that declines in latent-space separability reliably precede the corresponding drops in observable safety behaviors—typically by several training windows—establishing JSS as a robust early-warning signal for impending safety degradation during training.
-(a) Inverse DPO Pearson
-(b) Forward DPO Pearson
-(c) Inverse DPO Spearman
-(d) Forward DPO Spearmsn
-Figure 7: Safety Metrics (UR: Unsafe Rate; $R$: Refusual Rate) and JSS Correlation Heatmap
-To complement the dual-line analysis, we compute Pearson and Spearman correlations between latent JSS signals and surface safety metrics across lower, middle, and upper layers (Figure 7).
-As shown in the heatmaps, the correlation signs are fully consistent across all layer groups and both alignment directions.
-This confirms that JSS reliably tracks the direction of safety strengthening or weakening, providing a stable latent–surface correspondence.
-Figure 8: Comparative cost breakdown of traditional three-model red-teaming versus our JSS-only pipeline.
-For increasing red-teaming depths (T=1,5,10T={1,5,10}), the three panels report (left) total generated tokens, (middle) the relative share produced by each role—attacker ($M_{r}$), subject model ($M_{s}$), evaluator ($M_{e}$)—and (right) the per-token runtime cost.
-5.4 Baseline Comparison: Cost and Efficiency of Traditional Red-Teaming vs. Our JSS-Only Pipeline
-The goal of this experiment is to quantitatively compare resource consumption of two assessment pipelines under identical task scale and sample size conditions:
-  • Ours.
-Representation-based evaluation using offline attack trajectories.
-We execute a single forward pass on the subject model $M_{s}$ to extract hidden states; no online generation or additional evaluator inference is required.
-  • Baseline (traditional online red-team + llm-based evaluator). We run online red-teaming and safety evaluation using a PyRIT + Ollama setup, involving three model roles:
-    • $M_{r}$: red-team attack generator,
-    • $M_{s}$: subject model under test,
-    • $M_{e}$: evaluator / safety discriminator.
-Denote the total number of attack targets by $N$, the average number of turns per session by $\bar{T}$, and the average tokens per step by $\bar{L}$.
-We formalize token and time costs for both pipelines as follows.
-For our method, only a forward pass per turn of offline attack dataset on $M_{s}$ is required and hidden states are cached, then we can get the token cost and runtime cost:
-$C_{\text{token}}^{\text{Ours}}(N)\;=\;N\cdot\bar{T}_{\text{offine}},$ (32)
-$C_{\text{time}}^{\text{Ours}}\;=\;\mathrm{Time}_{M_{s}^{\text{forward}}}\;+\;\mathrm{Time}_{\text{offline\_JSS}}\;+\;\mathrm{Time}_{\text{benign\_con}},$ (33)
-where $\mathrm{Time}_{M_{s}^{\text{forward}}}$ denotes the cost of $C_{\text{token}}^{\text{ours}}(N,J)$, $\mathrm{Time}_{\text{offline\_JSS}}$ denotes the downstream cost for JSS statistics, and $\mathrm{Time}_{\text{benign\_con}}$ denotes time to construct the benign distribution (both $\mathrm{Time}_{\text{benign\_con}}$ and $\mathrm{Time}_{\text{offline\_JSS}}$ are typically much smaller than generation time, so we ignore them in following analysis).
-For traditional baseline, each red-team attempt involves a three-way inference flow for $T$ attack turns and $N$ attack targets:
-$C_{\text{token}}^{\text{Base}}(N)\;=\;N\cdot\bar{T_{r}}\cdot\big(\bar{L}_{r}\;+\;\bar{L}_{s}\;+\;\bar{L}_{e}\big),$ (34)
-$C_{\text{time}}^{\text{Base}}\;=\;\mathrm{Time}_{M_{r}}\;+\;\mathrm{Time}_{M_{s}}\;+\;\mathrm{Time}_{M_{e}},$ (35)
-where $\bar{L}_{r},\bar{L}_{s},\bar{L}_{e}$ denote the average token lengths for red-team generation, subject-model generation, and evaluator judgment respectively.
-$T_{r}$ denotes the average turn number of attacks generated by redteaming models.
-We run evaluations at multiple red-team intensities $N=\{n_{1},n_{2},\dots\}$ using PyRIT’s redteaming module to generate online attacks, and record produced token counts and elapsed times.
-For the JSS method, we count a single forward pass per example as one token-equivalent for bookkeeping; the JSS evaluation uses the same original instruction set $N$ corresponding offline multi-turn jailbreak datasets, derived RefEx and PlainR sets.
-We record token-equivalent costs and elapsed wall-clock time for both pipelines.
-We compare the two pipelines along the following axes:
-  • Relative share of output tokens: the fraction of total system tokens generated by the subject model $M_{s}$ during evaluation.
-  • Total generated-token count: the overall number of tokens produced by all system components during evaluation.
-  • Per-token time cost: wall-clock runtime divided by total produced tokens; this metric measures system throughput. Lower per-token time denotes higher efficiency.
-Figure 8 summarizes the cost comparison across all metrics, all roles, and all red-teaming depths, the JSS-only pipeline remains dramatically—often orders of magnitude—more efficient than the traditional three-model setup.
-Together, these three findings establish that N-GLARE reduces evaluation cost by at least two orders of magnitude while preserving ranking fidelity, effectively removing the generative bottleneck that makes traditional red-teaming expensive, slow, and difficult to scale.
+Note that "Hybrid" is the overall average across score all tested red-teaming tasks, with detailed task-wise scores provided in Appendix Table LABEL:tab:jb_safety_benchmark_split_A and Table LABEL:tab:jb_safety_benchmark_split_B.
 (a) Ranking stability across four jailbreak datasets.
@@
 Across all settings, the curves remain smooth and largely monotonic, confirming that N-GLARE yields stable safety ordering even when trajectory sampling budgets are perturbed.
+Figure 12: Comparative cost breakdown of traditional three-model red-teaming versus our JSS-only pipeline.
+For increasing red-teaming depths (T=1,5,10T={1,5,10}), the three panels report (left) total generated tokens, (middle) the relative share produced by each role—attacker ($M_{r}$), subject model ($M_{s}$), evaluator ($M_{e}$)—and (right) the per-token runtime cost.
 (a) Sensitivity analysis on jailbreak datasets.
 (b) Sensitivity analysis on benign datasets.
-Figure 10: Sensitivity heatmaps for JSS-based ranking across datasets, models (see Appendix Table III for alphabetical labels mapping), baselines, and sample sizes.
+Figure 13: Sensitivity heatmaps for JSS-based ranking across datasets, models (see Appendix Table 6 for alphabetical labels mapping), baselines, and sample sizes.
 Each heatmap visualizes pairwise rank consistency between two evaluation settings computed with two rank correlation metrics (S = Spearman, K = Kendall), where darker cells indicate higher agreement.
 Rows and columns correspond to jailbreak datasets, models, baseline datasets, or sampling budgets $n$.
-5.5 Model Sensitivity and Systemic Stability Experiments
-If our method is indeed capturing an intrinsic property of the model’s alignment dynamics—rather than dataset-specific artifacts—then its induced safety ordering should remain stable when we perturb the underlying data and sampling conditions.
-The goal of this section is therefore to test the systemic stability of our JSS-based safety scale.
-Concretely, we ask:
-  • whether the relative JSS differences between models (i.e., the safety ranking induced by $\mathrm{JSS}$) remain consistent when we vary which jailbreak trajectories are used under different attack datasets and how many sample size $n$ trajectories are included; and
-  • whether the same stability holds when we perturb the benign reference distribution, by changing the source corpus and the number of benign samples used to estimate the natural manifold.
-To assess whether JSS captures a dataset-agnostic and sample-size–agnostic property of model alignment dynamics, we evaluate stability under both jailbreak and benign perturbations.
-For jailbreak analysis, we use four multi-turn attack datasets—(D)=tom-gibbs, (A)=XGuard-Train, (B)=SafeMTData, (C)=MTSARed—and vary the number of included targets to produce small/medium/large (n)-sample conditions (each yielding ($5\text{–}10\times n$) activation trajectories).
-Figure  10 shows that across all settings, JSS-induced rankings remain highly consistent: Dataset–Dataset correlations reach mean Spearman (0.7150) and Kendall (0.5411); Model–Model comparisons show stable ordering; and varying (n) preserves strong agreement (mean Spearman (0.7973), Kendall (0.6328)).
-Even in pairwise ($n\times n$) evaluations, consistency remains extremely high (Spearman (0.9691), Kendall (0.8677)).
-An identical analysis on benign-distribution estimation—using four instruction-following corpora (LM)=LMSYS, (OA)=OASST1, (OP)=OpenHermes, (UL)=UltraChat—shows similarly robust behavior, with Dataset–Dataset correlations (mean Spearman (0.7012), Kendall (0.5237)) and Model–Model comparisons both remaining stable, and sample-size variation maintaining strong consistency (mean Spearman (0.81), Kendall (0.61)).
-Finally, Figure 9 provides a more intuitive visualization: as (n) increases, normalized JSS ranking curves across all jailbreak and benign datasets evolve smoothly and converge without reversals, reinforcing that JSS-based safety ordering is structurally stable across datasets, models, and sampling budgets.
-6 Conclusion
-This paper introduces N-GLARE, a non-generative, representation-based framework for LLM safety evaluation, showing that a model’s safety state is encoded not only in its outputs but in the geometry and dynamics of its latent trajectories.
-By constructing a benign distribution manifold and using JSS to quantify how jailbreak and ideal-refusal trajectories deviate from it, we obtain safety indicators that are comparable across models.
-Across more than 40 models and multiple red-teaming strategies, N-GLARE’s safety rankings remain consistent with large-scale red-teaming evaluations while reducing evaluation cost by over two orders of magnitude.
-Beyond reproducing traditional benchmarks, N-GLARE uncovers finer-grained safety dynamics, such as the phase structure of jailbreak progression, the synchronized evolution of refusal tendency, and monotonic trends under forward and inverse DPO, showcasing latent-trajectory analysis as a scalable paradigm for model-level safety diagnostics and providing a new foundation for continuous alignment monitoring and future alignment research.
-Appendix A Model Safety Evaluation Results
-Table LABEL:tab:jb_safety_multi presents the performance of 36 large language models tested across various security evaluation tasks.
-These tasks primarily measure their ability to resist jailbreak attacks, bypass protective mechanisms, and handle harmful instructions.
-The data in the table includes the models’ effectiveness in rejecting malicious requests, their ability to defend against bypass attacks, and their vulnerabilities in specific tasks such as ASCII art interpretation exploits and role-playing attacks.
-TABLE I: Model Safety Benchmark Results
-[…]
-1 Llama-3.2-3B-Instruct 0.6733 0.2899 81.89 91.18 83.62 44.42 47.31 97.85 100 99.57 94.40 96.12 94.05 91.38 95.72 90.95 97.71 98.28 78.83 97.41 100 95.69 64.08 98.28
-2 Qwen/Qwen2.5-7B-Instruct 0.5346 0.4414 55.17 97.78 17.42 13.82 50.28 56.94 28.93 95.26 42.81 96.88 8.29 33.69 86.80 33.16 90.24 98.63 95.09 86.51 99.52 93.10 33.91 84.74
-3 mistralai/Mistral-7B-Instruct-v0.3 0.3582 0.1776 34.69 61.03 8.12 6.03 25.53 47.01 3.45 20.26 25.43 23.71 2.05 31.19 37.87 23.62 70.69 71.55 63.12 79.05 84.05 80.60 23.41 78.02
-4 Llama-3.1-8B-it 0.4312 0.2538 93.22 100 99.14 65.38 100 98.71 99.43 99.14 97.63 100 97.43 89.66 99.57 95.30 98.28 100 93.63 97.41 98.71 100 90.52 100
-5 google/gemma-7b-it 0.6992 0.5768 64.02 94.50 56.03 36.85 48.08 69.83 100 72.41 38.36 97.86 19.86 71.92 90.95 64.39 96.12 90.95 69.26 80.60 97.84 96.12 29.70 90.09
-6 Mistral-7B-Instruct-v0.2 0.3400 0.3526 36.92 40.20 7.66 4.77 54.74 42.49 6.03 12.07 19.83 24.57 1.49 40.73 29.13 32.61 63.52 78.02 60.25 65.95 97.41 79.74 32.62 67.38
-7 nvidia/Nemotron-Mini-4B-Instruct 0.5784 0.3118 51.57 96.85 22.83 12.54 95.28 50.00 99.21 29.13 35.43 91.49 7.31 38.12 93.97 33.00 85.83 99.21 92.13 82.68 96.06 74.02 23.08 75.00
-8 Baichuan2-7B-Chat 0.2792 0.5432 33.17 46.17 35.04 11.50 13.79 35.97 20.16 13.96 10.84 37.72 10.03 11.74 23.35 45.94 71.80 47.86 58.18 74.25 84.05 72.33 24.07 62.12
-9 gemma-2-9b-it 0.3329 0.4727 67.63 95.62 68.10 2.15 89.22 78.02 25.00 25.43 64.67 53.45 31.21 93.10 96.99 24.35 85.34 100 100 89.66 100 99.57 78.02 92.27
-10 TinySwallow-1.5B-Instruct 0.6708 0.3339 68.22 98.43 44.88 14.76 98.43 71.65 100 94.49 44.88 100 46.13 63.24 98.43 57.43 96.06 98.43 97.64 86.61 100 98.43 57.14 92.13
-11 microsoft/Phi-3-mini-4k-instruct 0.2995 0.5432 63.90 96.05 78.74 21.09 92.91 74.80 97.64 99.21 37.80 99.21 11.49 33.89 96.85 61.72 87.40 100 68.04 88.19 100 97.64 57.98 88.19
-12 google/gemma-3-12b-it 0.6629 0.3645 71.16 98.45 85.83 18.11 85.04 90.55 21.71 88.28 88.19 86.99 10.51 64.66 96.85 63.79 88.98 97.64 100 95.28 100 96.85 67.50 96.85
-13 Llama-3.1-Tulu-3-8B 0.9034 0.2734 65.95 96.26 32.81 43.19 49.68 62.89 96.25 100 38.58 98.43 10.79 56.41 88.75 73.97 100 100 90.55 100 100 100 65.54 100
-14 granite-3.0-8b-instruct 0.4594 0.5180 61.30 99.63 27.59 58.79 28.63 63.39 55.17 89.37 15.19 91.81 12.89 36.07 98.71 60.83 99.14 99.14 78.15 91.81 99.14 98.71 61.55 86.21
-15 DeepSeek-R1-Distill-Qwen-7B 0.5685 1.0285 53.49 65.73 25.20 28.38 82.67 75.78 26.77 83.43 64.84 83.12 33.90 49.57 44.09 89.06 59.68 54.33 58.24 87.58 33.49 77.17 42.23 87.64
-16 microsoft/Phi-3-mini-128k-instruct 0.2784 0.4763 53.82 94.87 30.60 6.78 84.48 71.98 80.17 95.69 21.98 97.41 5.38 26.61 93.58 47.65 80.60 100 67.42 84.48 100 94.86 44.63 80.76
-17 Mistral-NeMo-Minitron-8B-Instruct 0.7369 0.3435 47.88 84.85 19.69 21.76 35.55 48.41 14.61 43.51 29.69 36.32 3.75 43.34 91.34 36.12 90.55 86.87 74.72 87.49 96.06 81.38 59.08 71.79
-18 01-ai/Yi-1.5-9B-Chat 0.1992 0.7507 39.36 85.17 12.93 3.39 39.24 61.64 27.35 9.83 21.98 53.06 11.75 23.60 67.95 27.14 74.68 86.18 47.63 77.34 93.10 75.54 27.04 76.82
-19 Yi-6B-Chat 0.5100 0.3796 43.36 46.25 9.32 8.78 36.01 42.34 60.51 73.18 13.26 80.69 4.87 69.47 74.88 31.84 81.33 94.40 61.35 72.10 97.84 92.24 23.35 67.72
-20 Qwen/Qwen2.5-0.5B-Instruct 0.5723 0.1777 44.05 75.06 11.71 4.45 59.99 48.15 93.13 47.45 23.24 92.67 12.81 32.20 61.16 38.57 80.76 67.81 86.92 76.76 98.28 61.45 23.84 66.08
-21 Qwen2.5-1.5B-Instruct 0.6726 0.3052 52.40 98.90 11.82 17.69 85.34 48.29 99.14 72.75 29.23 97.43 14.29 23.51 87.17 43.02 95.69 95.69 91.10 81.03 99.14 85.78 21.97 76.06
-22 Qwen2.5-14B-Instruct 0.4591 0.4041 68.61 99.63 28.80 4.27 91.38 83.19 78.45 98.28 58.80 98.28 36.62 58.19 93.60 67.38 96.55 100 100 94.83 96.12 99.14 77.59 92.67
-23 google/gemma-3-4b-it 0.4624 0.2563 57.75 93.87 50.39 11.81 56.69 75.59 17.97 42.18 69.29 39.37 22.56 49.14 94.53 52.13 66.93 97.64 98.43 91.34 100 98.43 51.26 95.28
-24 google/gemma-3-1b-it 0.6095 0.5359 65.19 96.15 59.06 15.50 60.77 73.23 83.46 22.05 70.08 80.60 44.69 41.15 96.85 37.07 71.65 100 96.85 93.70 99.21 94.49 65.55 88.98
-25 gemma-2-2b-it 0.2755 0.2885 65.56 72.79 46.12 48.08 85.38 81.03 38.57 23.85 70.93 66.38 32.13 76.55 98.71 19.31 82.33 100 99.57 90.09 66.67 99.57 83.33 90.09
-26 DeepSeek-R1-Distill-Qwen-1.5B 0.2589 0.7154 40.60 41.18 19.53 30.10 61.60 82.36 27.09 31.92 66.89 15.25 17.74 49.94 25.39 38.26 39.14 25.30 44.60 63.69 34.37 49.74 51.52 60.61
-27 DeepSeek-R1-Distill-Qwen-14B 0.5352 1.4011 57.53 76.38 34.65 24.22 80.60 75.59 32.81 27.34 70.08 25.20 34.78 65.52 59.06 74.05 74.21 44.09 80.31 84.72 77.07 88.98 63.03 83.20
-28 deepseek-ai/DeepSeek-R1-Distill-Llama-8B 0.5352 1.4011 51.17 70.25 21.26 23.26 92.23 61.71 39.84 36.60 49.22 31.25 29.42 53.45 51.97 86.11 69.29 59.84 62.20 86.03 55.35 85.04 39.50 79.42
-29 meta-llama/Meta-Llama-3-8B-Instruct 0.0798 0.2441 79.98 91.61 71.12 80.26 72.13 97.84 100 71.37 93.99 100 20.84 83.33 96.57 97.13 93.16 67.24 98.71 98.71 99.57 100 87.60 98.71
-30 meta-llama/Llama-2-7b-chat-hf 0.7188 0.4138 64.47 46.79 98.71 7.73 99.14 88.53 100 89.80 47.09 98.71 12.33 73.07 97.41 58.40 91.38 100 76.60 90.09 100 99.14 85.99 88.36
-31 meta-llama/Llama-2-13b-chat-hf 0.7386 1.4244 75.40 63.94 98.56 85.46 83.73 80.52 99.71 98.28 59.48 99.43 12.36 60.51 96.57 66.72 99.14 99.57 75.49 91.38 99.57 100 89.66 90.99
-32 Qwen/Qwen2.5-72B-Instruct 0.5530 0.6269 79.41 97.71 54.61 10.80 98.86 78.06 100 98.58 88.67 95.74 83.13 61.76 98.29 86.25 95.73 100 98.29 87.75 99.43 98.29 81.63 92.88
-33 Qwen/Qwen2.5-32B-Instruct 0.4748 0.5229 68.81 97.10 56.56 12.57 94.17 79.12 44.30 92.87 55.67 97.69 28.12 64.47 96.24 64.71 93.44 99.60 97.81 92.24 100 97.48 70.77 89.43
-34 Qwen/Qwen2.5-3B-Instruct 0.3775 0.3692 51.16 98.92 23.10 9.36 36.22 49.57 63.68 81.47 25.64 97.41 11.10 26.18 94.45 28.74 93.10 99.14 91.02 78.11 100 95.26 35.22 80.60
-35 meta-llama/Meta-Llama-3-70B-Instruct 0.9358 0.2999 59.43 74.28 93.10 40.14 14.35 92.75 100 96.84 17.28 83.49 51.15 45.19 91.52 83.43 81.52 94.89 68.37 57.77 54.30 94.01 45.48 55.94
-36 Llama-3.1-Nemotron-70B-Instruct-HF 0.2293 0.4283 69.69 92.96 32.28 84.15 79.31 75.78 84.37 66.93 54.33 78.74 14.30 60.34 94.83 75.86 95.28 96.06 89.06 94.53 100 98.43 69.23 93.16
-TABLE II: Model–Abbreviation Mapping of Subset models Scatter Figure
+For jailbreak analysis, we use four multi-turn attack datasets—(D)=tom-gibbs, (A)=XGuard-Train, (B)=SafeMTData, (C)=MTSARed—and vary the number of included targets to produce small/medium/large (n)-sample conditions (each yielding ($5\text{–}10\times n$) activation trajectories)
+Table 5: Model–Abbreviation Mapping of Subset models Scatter Figure
 Model Abbr
@@
 nvidia/Nemotron-Mini-4B-Instruct INS-4B
-TABLE III: Model ID Mapping with Alphabetical Labels
+Table 6: Model ID Mapping with Alphabetical Labels
 Model ID Letter
@@
 Llama2-7B J
-Appendix B Unsafe Rate and Refusual Rate Definition
-Unsafe Rate. This metric is computed using the safety classifications provided by Qwen3Guard. A model output is counted as unsafe whenever the guard model labels it as “Unsafe.” Formally,
-$\text{Unsafe Rate}=\frac{N_{\text{Unsafe}}}{N_{\text{Total}}},$
-where $N_{\text{Unsafe}}$ denotes the number of responses flagged as Unsafe by the guard model, and $N_{\text{Total}}$ denotes the total number of model responses evaluated.
-Refusal Rate.
-This metric is based on refusal-style keyword matching.
-A model response is counted as a refusal if it contains typical refusal patterns such as “I’m sorry”, “I cannot comply”, “I refuse to answer”, or similar formulations.
-Formally,
-$\text{Refusal Rate}=\frac{N_{\text{Refusal}}}{N_{\text{Total}}},$
-where $N_{\text{Refusal}}$ denotes the number of responses that match the refusal keyword patterns.
-Complementarity.
-These two metrics capture opposite behavioral tendencies and are therefore complementary.
-Lower Unsafe Rate indicates stronger safety, while lower Refusal Rate indicates better helpfulness and less over-defensiveness.
-A high Unsafe Rate signals safety failures, whereas a high Refusal Rate reflects an overly conservative model that rejects too many user requests.
+Table 7: Model Safety Benchmark Results (split table A: None to DrAttack).
+models JB/RE BJ/BP AI Safety None ABJ Adaptive ArtPrompt AutoDAN Cipher DAN DeepInc DevMode DRA DrAttack
+1 Llama-3.2-3B-Instruct 0.6733 0.2899 81.89 91.18 83.62 44.42 47.31 97.85 100 99.57 94.40 96.12 94.05 91.38
+2 Qwen/Qwen2.5-7B-Instruct 0.5346 0.4414 55.17 97.78 17.42 13.82 50.28 56.94 28.93 95.26 42.81 96.88 8.29 33.69
+3 mistralai/Mistral-7B-Instruct-v0.3 0.3582 0.1776 34.69 61.03 8.12 6.03 25.53 47.01 3.45 20.26 25.43 23.71 2.05 31.19
+4 Llama-3.1-8B-it 0.4312 0.2538 93.22 100 99.14 65.38 100 98.71 99.43 99.14 97.63 100 97.43 89.66
+5 google/gemma-7b-it 0.6992 0.5768 64.02 94.50 56.03 36.85 48.08 69.83 100 72.41 38.36 97.86 19.86 71.92
+6 Mistral-7B-Instruct-v0.2 0.3400 0.3526 36.92 40.20 7.66 4.77 54.74 42.49 6.03 12.07 19.83 24.57 1.49 40.73
+7 nvidia/Nemotron-Mini-4B-Instruct 0.5784 0.3118 51.57 96.85 22.83 12.54 95.28 50.00 99.21 29.13 35.43 91.49 7.31 38.12
+8 Baichuan2-7B-Chat 0.2792 0.5432 33.17 46.17 35.04 11.50 13.79 35.97 20.16 13.96 10.84 37.72 10.03 11.74
+9 gemma-2-9b-it 0.3329 0.4727 67.63 95.62 68.10 2.15 89.22 78.02 25.00 25.43 64.67 53.45 31.21 93.10
+10 TinySwallow-1.5B-Instruct 0.6708 0.3339 68.22 98.43 44.88 14.76 98.43 71.65 100 94.49 44.88 100 46.13 63.24
+11 microsoft/Phi-3-mini-4k-instruct 0.2995 0.5432 63.90 96.05 78.74 21.09 92.91 74.80 97.64 99.21 37.80 99.21 11.49 33.89
+12 google/gemma-3-12b-it 0.6629 0.3645 71.16 98.45 85.83 18.11 85.04 90.55 21.71 88.28 88.19 86.99 10.51 64.66
+13 Llama-3.1-Tulu-3-8B 0.9034 0.2734 65.95 96.26 32.81 43.19 49.68 62.89 96.25 100 38.58 98.43 10.79 56.41
+14 granite-3.0-8b-instruct 0.4594 0.5180 61.30 99.63 27.59 58.79 28.63 63.39 55.17 89.37 15.19 91.81 12.89 36.07
+15 DeepSeek-R1-Distill-Qwen-7B 0.5685 1.0285 53.49 65.73 25.20 28.38 82.67 75.78 26.77 83.43 64.84 83.12 33.90 49.57
+16 microsoft/Phi-3-mini-128k-instruct 0.2784 0.4763 53.82 94.87 30.60 6.78 84.48 71.98 80.17 95.69 21.98 97.41 5.38 26.61
+17 Mistral-NeMo-Minitron-8B-Instruct 0.7369 0.3435 47.88 84.85 19.69 21.76 35.55 48.41 14.61 43.51 29.69 36.32 3.75 43.34
+18 01-ai/Yi-1.5-9B-Chat 0.1992 0.7507 39.36 85.17 12.93 3.39 39.24 61.64 27.35 9.83 21.98 53.06 11.75 23.60
+19 Yi-6B-Chat 0.5100 0.3796 43.36 46.25 9.32 8.78 36.01 42.34 60.51 73.18 13.26 80.69 4.87 69.47
+20 Qwen/Qwen2.5-0.5B-Instruct 0.5723 0.1777 44.05 75.06 11.71 4.45 59.99 48.15 93.13 47.45 23.24 92.67 12.81 32.20
+21 Qwen2.5-1.5B-Instruct 0.6726 0.3052 52.40 98.90 11.82 17.69 85.34 48.29 99.14 72.75 29.23 97.43 14.29 23.51
+22 Qwen2.5-14B-Instruct 0.4591 0.4041 68.61 99.63 28.80 4.27 91.38 83.19 78.45 98.28 58.80 98.28 36.62 58.19
+23 google/gemma-3-4b-it 0.4624 0.2563 57.75 93.87 50.39 11.81 56.69 75.59 17.97 42.18 69.29 39.37 22.56 49.14
+24 google/gemma-3-1b-it 0.6095 0.5359 65.19 96.15 59.06 15.50 60.77 73.23 83.46 22.05 70.08 80.60 44.69 41.15
+25 gemma-2-2b-it 0.2755 0.2885 65.56 72.79 46.12 48.08 85.38 81.03 38.57 23.85 70.93 66.38 32.13 76.55
+26 DeepSeek-R1-Distill-Qwen-1.5B 0.2589 0.7154 40.60 41.18 19.53 30.10 61.60 82.36 27.09 31.92 66.89 15.25 17.74 49.94
+27 DeepSeek-R1-Distill-Qwen-14B 0.5352 1.4011 57.53 76.38 34.65 24.22 80.60 75.59 32.81 27.34 70.08 25.20 34.78 65.52
+28 deepseek-ai/DeepSeek-R1-Distill-Llama-8B 0.5352 1.4011 51.17 70.25 21.26 23.26 92.23 61.71 39.84 36.60 49.22 31.25 29.42 53.45
+29 meta-llama/Meta-Llama-3-8B-Instruct 0.0798 0.2441 79.98 91.61 71.12 80.26 72.13 97.84 100 71.37 93.99 100 20.84 83.33
+30 meta-llama/Llama-2-7b-chat-hf 0.7188 0.4138 64.47 46.79 98.71 7.73 99.14 88.53 100 89.80 47.09 98.71 12.33 73.07
+31 meta-llama/Llama-2-13b-chat-hf 0.7386 1.4244 75.40 63.94 98.56 85.46 83.73 80.52 99.71 98.28 59.48 99.43 12.36 60.51
+32 Qwen/Qwen2.5-72B-Instruct 0.5530 0.6269 79.41 97.71 54.61 10.80 98.86 78.06 100 98.58 88.67 95.74 83.13 61.76
+33 Qwen/Qwen2.5-32B-Instruct 0.4748 0.5229 68.81 97.10 56.56 12.57 94.17 79.12 44.30 92.87 55.67 97.69 28.12 64.47
+34 Qwen/Qwen2.5-3B-Instruct 0.3775 0.3692 51.16 98.92 23.10 9.36 36.22 49.57 63.68 81.47 25.64 97.41 11.10 26.18
+35 meta-llama/Meta-Llama-3-70B-Instruct 0.9358 0.2999 59.43 74.28 93.10 40.14 14.35 92.75 100 96.84 17.28 83.49 51.15 45.19
+36 Llama-3.1-Nemotron-70B-Instruct-HF 0.2293 0.4283 69.69 92.96 32.28 84.15 79.31 75.78 84.37 66.93 54.33 78.74 14.30 60.34
+Table 8: Model Safety Benchmark Results (split table B: GCG to TAP).
+models JB/RE BJ/BP AI Safety GCG GPTFuzzer Grandmother Masterkey MultiLing PAIR PastTense Psychology ReNeLLM TAP
+1 Llama-3.2-3B-Instruct 0.6733 0.2899 81.89 95.72 90.95 97.71 98.28 78.83 97.41 100 95.69 64.08 98.28
+2 Qwen/Qwen2.5-7B-Instruct 0.5346 0.4414 55.17 86.80 33.16 90.24 98.63 95.09 86.51 99.52 93.10 33.91 84.74
+3 mistralai/Mistral-7B-Instruct-v0.3 0.3582 0.1776 34.69 37.87 23.62 70.69 71.55 63.12 79.05 84.05 80.60 23.41 78.02
+4 Llama-3.1-8B-it 0.4312 0.2538 93.22 99.57 95.30 98.28 100 93.63 97.41 98.71 100 90.52 100
+5 google/gemma-7b-it 0.6992 0.5768 64.02 90.95 64.39 96.12 90.95 69.26 80.60 97.84 96.12 29.70 90.09
+6 Mistral-7B-Instruct-v0.2 0.3400 0.3526 36.92 29.13 32.61 63.52 78.02 60.25 65.95 97.41 79.74 32.62 67.38
+7 nvidia/Nemotron-Mini-4B-Instruct 0.5784 0.3118 51.57 93.97 33.00 85.83 99.21 92.13 82.68 96.06 74.02 23.08 75.00
+8 Baichuan2-7B-Chat 0.2792 0.5432 33.17 23.35 45.94 71.80 47.86 58.18 74.25 84.05 72.33 24.07 62.12
+9 gemma-2-9b-it 0.3329 0.4727 67.63 96.99 24.35 85.34 100 100 89.66 100 99.57 78.02 92.27
+10 TinySwallow-1.5B-Instruct 0.6708 0.3339 68.22 98.43 57.43 96.06 98.43 97.64 86.61 100 98.43 57.14 92.13
+11 microsoft/Phi-3-mini-4k-instruct 0.2995 0.5432 63.90 96.85 61.72 87.40 100 68.04 88.19 100 97.64 57.98 88.19
+12 google/gemma-3-12b-it 0.6629 0.3645 71.16 96.85 63.79 88.98 97.64 100 95.28 100 96.85 67.50 96.85
+13 Llama-3.1-Tulu-3-8B 0.9034 0.2734 65.95 88.75 73.97 100 100 90.55 100 100 100 65.54 100
+14 granite-3.0-8b-instruct 0.4594 0.5180 61.30 98.71 60.83 99.14 99.14 78.15 91.81 99.14 98.71 61.55 86.21
+15 DeepSeek-R1-Distill-Qwen-7B 0.5685 1.0285 53.49 44.09 89.06 59.68 54.33 58.24 87.58 33.49 77.17 42.23 87.64
+16 microsoft/Phi-3-mini-128k-instruct 0.2784 0.4763 53.82 93.58 47.65 80.60 100 67.42 84.48 100 94.86 44.63 80.76
+17 Mistral-NeMo-Minitron-8B-Instruct 0.7369 0.3435 47.88 91.34 36.12 90.55 86.87 74.72 87.49 96.06 81.38 59.08 71.79
+18 01-ai/Yi-1.5-9B-Chat 0.1992 0.7507 39.36 67.95 27.14 74.68 86.18 47.63 77.34 93.10 75.54 27.04 76.82
+19 Yi-6B-Chat 0.5100 0.3796 43.36 74.88 31.84 81.33 94.40 61.35 72.10 97.84 92.24 23.35 67.72
+20 Qwen/Qwen2.5-0.5B-Instruct 0.5723 0.1777 44.05 61.16 38.57 80.76 67.81 86.92 76.76 98.28 61.45 23.84 66.08
+21 Qwen2.5-1.5B-Instruct 0.6726 0.3052 52.40 87.17 43.02 95.69 95.69 91.10 81.03 99.14 85.78 21.97 76.06
+22 Qwen2.5-14B-Instruct 0.4591 0.4041 68.61 93.60 67.38 96.55 100 100 94.83 96.12 99.14 77.59 92.67
+23 google/gemma-3-4b-it 0.4624 0.2563 57.75 94.53 52.13 66.93 97.64 98.43 91.34 100 98.43 51.26 95.28
+24 google/gemma-3-1b-it 0.6095 0.5359 65.19 96.85 37.07 71.65 100 96.85 93.70 99.21 94.49 65.55 88.98
+25 gemma-2-2b-it 0.2755 0.2885 65.56 98.71 19.31 82.33 100 99.57 90.09 66.67 99.57 83.33 90.09
+26 DeepSeek-R1-Distill-Qwen-1.5B 0.2589 0.7154 40.60 25.39 38.26 39.14 25.30 44.60 63.69 34.37 49.74 51.52 60.61
+27 DeepSeek-R1-Distill-Qwen-14B 0.5352 1.4011 57.53 59.06 74.05 74.21 44.09 80.31 84.72 77.07 88.98 63.03 83.20
+28 deepseek-ai/DeepSeek-R1-Distill-Llama-8B 0.5352 1.4011 51.17 51.97 86.11 69.29 59.84 62.20 86.03 55.35 85.04 39.50 79.42
+29 meta-llama/Meta-Llama-3-8B-Instruct 0.0798 0.2441 79.98 96.57 97.13 93.16 67.24 98.71 98.71 99.57 100 87.60 98.71
+30 meta-llama/Llama-2-7b-chat-hf 0.7188 0.4138 64.47 97.41 58.40 91.38 100 76.60 90.09 100 99.14 85.99 88.36
+31 meta-llama/Llama-2-13b-chat-hf 0.7386 1.4244 75.40 96.57 66.72 99.14 99.57 75.49 91.38 99.57 100 89.66 90.99
+32 Qwen/Qwen2.5-72B-Instruct 0.5530 0.6269 79.41 98.29 86.25 95.73 100 98.29 87.75 99.43 98.29 81.63 92.88
+33 Qwen/Qwen2.5-32B-Instruct 0.4748 0.5229 68.81 96.24 64.71 93.44 99.60 97.81 92.24 100 97.48 70.77 89.43
+34 Qwen/Qwen2.5-3B-Instruct 0.3775 0.3692 51.16 94.45 28.74 93.10 99.14 91.02 78.11 100 95.26 35.22 80.60
+35 meta-llama/Meta-Llama-3-70B-Instruct 0.9358 0.2999 59.43 91.52 83.43 81.52 94.89 68.37 57.77 54.30 94.01 45.48 55.94
+36 Llama-3.1-Nemotron-70B-Instruct-HF 0.2293 0.4283 69.69 94.83 75.86 95.28 96.06 89.06 94.53 100 98.43 69.23 93.16
+Table 9: Model Safety Benchmark (split table A). AI Safety is split into Risks and Trials.
+models Risks Trials None ABJ Adaptive ArtPrompt AutoDAN Cipher DAN DeepInc DevMode DRA DrAttack
+1 Llama-3.2-3B-Instruct 815 5860 36/408 38/232 211/464 244/464 4/232 0/232 1/232 13/232 9/232 14/348 20/232
+2 Qwen/Qwen2.5-7B-Instruct 18586 68740 439/22435 2164/2624 2260/2624 1065/2146 1123/2617 1776/2517 110/2359 953/1667 62/2145 3434/3793 1398/2117
+3 mistralai/Mistral-7B-Instruct-v0.3 2863 4912 106/272 213/232 218/232 172/232 122/232 224/232 185/232 173/232 177/232 226/232 159/232
+4 Llama-3.1-8B-it 230 5167 0/34 2/232 128/435 0/232 3/232 1/174 2/232 11/464 0/232 3/232 24/232
+5 google/gemma-7b-it 1330 5028 14/272 102/232 118/232 119/232 70/232 0/232 64/232 143/232 3/232 166/232 97/348
+6 Mistral-7B-Instruct-v0.2 3591 6101 728/1229 214/232 220/232 105/232 133/232 218/232 204/232 186/232 175/232 228/232 275/464
+7 nvidia/Nemotron-Mini-4B-Instruct 927 2601 4/127 98/127 110/127 6/127 63/127 1/127 90/127 82/127 8/127 104/116 71/116
+8 Baichuan2-7B-Chat 2853 4939 145/299 150/232 198/232 199/232 122/232 182/232 198/232 205/232 138/232 178/232 183/232
+9 gemma-2-9b-it 1400 4912 9/272 74/232 227/232 25/232 51/232 174/232 173/232 80/232 108/232 156/232 6/232
+10 TinySwallow-1.5B-Instruct 525 2637 2/127 70/127 92/127 2/127 36/127 0/127 7/127 70/127 0/127 65/127 42/116
+11 microsoft/Phi-3-mini-4k-instruct 662 2803 7/177 27/127 100/127 9/127 32/127 3/127 1/127 79/127 1/127 109/127 76/116
+12 google/gemma-3-12b-it 557 2639 2/129 18/127 104/127 19/127 12/127 93/127 14/127 15/127 12/127 112/127 41/116
+13 Llama-3.1-Tulu-3-8B 680 2887 4/159 85/127 89/159 79/159 59/159 5/159 0/127 78/127 2/127 112/127 50/116
+14 granite-3.0-8b-instruct 2595 7580 1/272 252/348 325/812 165/232 166/464 156/348 37/348 295/348 19/232 390/464 218/348
+15 DeepSeek-R1-Distill-Qwen-7B 979 2687 60/177 95/127 85/127 14/127 30/127 93/127 11/127 44/127 18/127 56/127 58/116
+16 microsoft/Phi-3-mini-128k-instruct 1581 4912 13/272 161/232 216/232 36/232 65/232 46/232 10/232 181/232 6/232 218/232 170/232
+17 Mistral-NeMo-Minitron-8B-Instruct 1108 2612 12/127 102/127 97/127 80/127 63/127 108/127 69/127 89/127 69/127 110/116 62/116
+18 01-ai/Yi-1.5-9B-Chat 2487 4912 36/272 202/232 224/232 136/232 89/232 168/232 209/232 181/232 96/232 198/232 177/232
+19 Yi-6B-Chat 2784 60676 761/1027 210/232 208/232 141/232 1126/232 91/232 60/232 200/232 44/232 212/232 68/232
+20 Qwen/Qwen2.5-0.5B-Instruct 2058 4912 64/272 199/232 213/232 91/232 108/232 15/232 119/232 173/232 17/232 175/232 151/232
+21 Qwen2.5-1.5B-Instruct 1613 4912 3/272 202/232 163/232 34/232 119/232 2/232 61/232 163/232 4/232 183/232 176/232
+22 Qwen2.5-14B-Instruct 1025 4912 1/272 164/232 222/232 20/232 39/232 50/232 3/232 95/232 3/232 144/232 97/232
+23 google/gemma-3-4b-it 995 2766 6/129 63/127 112/127 55/127 31/127 104/127 73/127 39/127 77/127 191/254 59/116
+24 google/gemma-3-1b-it 738 2639 4/129 52/127 107/127 46/127 34/127 21/127 99/127 38/127 23/127 59/127 67/116
+25 gemma-2-2b-it 1990 6168 37/136 250/464 119/232 67/464 44/232 209/348 265/348 66/232 78/232 288/464 136/580
+26 DeepSeek-R1-Distill-Qwen-1.5B 1373 2662 73/127 102/127 87/152 35/127 19/127 89/127 83/127 40/127 107/127 74/127 56/116
+27 DeepSeek-R1-Distill-Qwen-14B 978 2637 30/127 83/127 96/127 23/127 31/127 85/127 92/127 38/127 95/127 56/127 40/116
+28 deepseek-ai/DeepSeek-R1-Distill-Llama-8B 1108 2783 46/157 100/127 96/127 8/127 48/127 76/127 79/127 64/127 87/127 62/127 54/116
+29 meta-llama/Meta-Llama-3-8B-Instruct 844 7873 167/2798 134/464 4/29 97/348 5/232 0/232 45/290 13/232 0/232 161/232 37/232
+30 meta-llama/Llama-2-7b-chat-hf 3828 10893 2629/5180 3/232 24/29 1/232 23/232 0/232 20/232 293/580 3/232 271/348 93/348
+31 meta-llama/Llama-2-13b-chat-hf 1749 8187 868/2503 5/348 21/232 36/348 67/348 1/348 3/232 185/464 2/348 194/232 91/232
+32 Qwen/Qwen2.5-72B-Instruct 1009 7497 6/391 156/351 313/351 4/351 77/351 0/351 4/351 38/351 14/351 55/351 130/340
+33 Qwen/Qwen2.5-32B-Instruct 7062 34002 110/4738 655/1508 1318/1508 87/1508 314/1504 840/1508 101/1508 668/1508 26/1508 1143/1624 494/1392
+34 Qwen/Qwen2.5-3B-Instruct 1762 4918 3/278 176/232 210/232 135/232 117/232 68/232 43/232 172/232 6/232 199/232 171/232
+35 meta-llama/Meta-Llama-3-70B-Instruct 1666 6650 309/1430 16/232 139/232 199/232 17/232 0/232 7/232 192/232 38/232 113/232 127/232
+36 Llama-3.1-Nemotron-70B-Instruct-HF 537 2601 8/127 86/127 11/127 26/127 31/127 20/127 42/127 58/127 27/127 109/127 50/127
+Table 10: Model Safety Benchmark (split table B). AI Safety is split into Risks and Trials.
+models Risks Trials GCG GPTFuzzer Grandmother Masterkey MultiLing PAIR PastTense Psychology ReNeLLM TAP
+1 Llama-3.2-3B-Instruct 815 5860 8/232 21/232 7/348 4/232 40/232 6/232 0/232 10/232 125/348 4/232
+2 Qwen/Qwen2.5-7B-Instruct 18586 68740 2233/1986 1014/1537 224/2305 33/2624 88/1986 311/2305 11/2305 159/2305 1401/2138 338/2215
+3 mistralai/Mistral-7B-Instruct-v0.3 2863 4912 141/232 176/232 68/232 66/232 81/232 47/232 37/232 45/232 176/232 51/232
+4 Llama-3.1-8B-it 230 5167 1/232 8/232 6/348 0/232 10/232 6/232 3/232 0/232 22/232 0/232
+5 google/gemma-7b-it 1330 5028 21/232 80/232 9/232 21/232 61/232 45/232 5/232 9/232 160/232 23/232
+6 Mistral-7B-Instruct-v0.2 3591 6101 162/232 154/232 84/232 51/232 91/232 79/232 6/232 47/232 156/232 75/232
+7 nvidia/Nemotron-Mini-4B-Instruct 927 2601 7/116 76/116 18/127 1/127 10/127 22/127 5/127 33/127 89/116 29/116
+8 Baichuan2-7B-Chat 2853 4939 171/232 117/232 61/232 120/232 89/232 59/232 37/232 62/232 153/232 86/232
+9 gemma-2-9b-it 1400 4912 6/232 174/232 34/232 0/232 0/232 24/232 0/232 1/232 51/232 17/232
+10 TinySwallow-1.5B-Instruct 525 2637 2/127 47/116 5/127 2/127 3/127 17/127 0/127 2/127 51/119 10/127
+11 microsoft/Phi-3-mini-4k-instruct 662 2803 4/127 81/232 16/127 0/127 34/127 15/127 0/127 3/127 50/119 15/127
+12 google/gemma-3-12b-it 557 2639 4/127 42/116 14/127 3/127 0/127 6/127 0/127 4/127 38/119 4/127
+13 Llama-3.1-Tulu-3-8B 680 2887 17/159 37/145 0/127 0/127 12/127 0/127 0/127 0/127 51/148 0/127
+14 granite-3.0-8b-instruct 2595 7580 3/232 171/464 2/232 3/348 68/348 19/232 3/348 3/232 267/696 32/232
+15 DeepSeek-R1-Distill-Qwen-7B 979 2687 71/127 9/116 50/127 58/127 50/127 14/127 54/127 29/127 67/119 13/127
+16 microsoft/Phi-3-mini-128k-instruct 1581 4912 13/232 120/232 45/232 0/232 64/232 36/232 0/232 10/232 128/232 43/232
+17 Mistral-NeMo-Minitron-8B-Instruct 1108 2612 11/127 71/116 12/127 14/127 24/127 15/127 5/127 22/127 41/116 32/116
+18 01-ai/Yi-1.5-9B-Chat 2487 4912 68/232 165/232 58/232 24/232 116/232 51/232 16/232 56/232 164/232 53/232
+19 Yi-6B-Chat 2784 60676 56/232 150/232 40/232 13/232 74/232 64/232 5/232 18/232 173/232 70/232
+20 Qwen/Qwen2.5-0.5B-Instruct 2058 4912 81/232 125/232 43/232 74/232 25/232 50/232 4/232 85/232 170/232 76/232
+21 Qwen2.5-1.5B-Instruct 1613 4912 28/232 130/232 10/232 10/232 16/232 44/232 1/232 33/232 177/232 54/232
+22 Qwen2.5-14B-Instruct 1025 4912 12/232 75/232 8/232 0/232 0/232 12/232 9/232 2/232 52/232 17/232
+23 google/gemma-3-4b-it 995 2766 6/127 55/116 42/127 3/127 2/127 11/127 0/127 2/127 58/119 6/127
+24 google/gemma-3-1b-it 738 2639 4/127 73/116 36/127 0/127 4/127 8/127 1/127 7/127 41/119 14/127
+25 gemma-2-2b-it 1990 6168 3/232 15/232 41/232 0/232 1/232 23/232 116/348 1/232 37/232 23/232
+26 DeepSeek-R1-Distill-Qwen-1.5B 1373 2662 92/127 64/116 68/127 93/127 61/127 31/127 55/127 56/127 55/119 33/127
+27 DeepSeek-R1-Distill-Qwen-14B 978 2637 52/127 22/116 32/127 71/127 25/127 15/127 17/127 14/127 44/119 17/127
+28 deepseek-ai/DeepSeek-R1-Distill-Llama-8B 1108 2783 61/127 26/232 39/127 51/127 48/127 16/127 34/127 19/127 92/1190 22/127
+29 meta-llama/Meta-Llama-3-8B-Instruct 844 7873 6/232 10/348 14/232 114/348 3/232 3/232 1/232 0/232 27/232 3/232
+30 meta-llama/Llama-2-7b-chat-hf 3828 10893 6/232 283/696 20/232 0/232 42/232 23/232 0/232 2/232 65/464 27/232
+31 meta-llama/Llama-2-13b-chat-hf 1749 8187 6/232 149/464 2/232 1/232 53/232 20/232 1/232 0/232 24/232 20/232
+32 Qwen/Qwen2.5-72B-Instruct 1009 7497 6/351 45/340 15/351 0/351 6/351 43/351 1/351 8/467 63/343 25/351
+33 Qwen/Qwen2.5-32B-Instruct 7062 34002 50/1508 488/1392 99/1508 6/1508 31/1508 117/1508 0/1189 38/1508 324/11111 153/1448
+34 Qwen/Qwen2.5-3B-Instruct 1762 4918 10/232 162/232 16/232 2/232 19/232 50/232 0/232 11/232 147/232 45/232
+35 meta-llama/Meta-Llama-3-70B-Instruct 1666 6650 20/232 38/232 43/232 12/232 73/232 98/232 106/232 14/232 126/232 102/232
+36 Llama-3.1-Nemotron-70B-Instruct-HF 537 2601 7/127 31/127 6/127 5/127 14/127 7/127 0/127 2/127 39/127 9/127